### v2

# The geometry and arithmetic of bielliptic Picard curves

Abstract.
We study the geometry and arithmetic of the curves $C\colon y^{3}=x^{4}+ax^{2}+b$ and their associated Prym abelian surfaces $P$. We prove a Torelli theorem in this context and give a geometric proof of the fact that $P$ has quaternionic multiplication (QM) by the quaternion order of discriminant $6$. This allows us to describe the Galois action on the geometric endomorphism algebra of $P$. As an application, we classify the torsion subgroups of the Mordell-Weil groups $P({\rm\mathbb{Q}})$, as both abelian groups and ${\rm End}(P)$-modules.
1. Introduction
Let $k$ be a field of characteristic neither $2$ nor $3$. A bielliptic Picard curve over $k$ is a smooth projective curve $C$ with an affine model of the form
(1.1) $y^{3}=x^{4}+ax^{2}+b$
for some $a,b\in k$. Such a curve is equipped with both a $\mu_{3}$-action $(x,y)\mapsto(x,\omega y)$ and a commuting involution $\tau\colon(x,y)\mapsto(-x,y)$. The induced involution $\tau^{*}$ on the Jacobian variety $J=\operatorname{Jac}_{C}$ allows us to define the Prym variety $P=\ker(1+\tau^{*})$. The abelian surface $P/k$ inherits a $\mu_{3}$-action and carries a $(1,2)$-polarization; it need not be principally polarizable over $k$.
The goal of this paper is to explore the remarkably rich geometry and arithmetic of bielliptic Picard curves and their associated Pryms.
1.1. Basic results
We first prove some foundational results for bielliptic Picard curves. We
  1. (1)
    prove a Torelli theorem for the association $C\mapsto P$, once suitable data on both sides is fixed (Theorem 3.13);
  2. (2)
    show ${\rm End}(P_{\bar{k}})$ contains a maximal order $\mathcal{O}$ in a discriminant $6$ quaternion algebra (Proposition 6.1), in other words $P_{\bar{k}}$ has quaternionic multiplication (QM) by $\mathcal{O}$;
  3. (3)
    give an explicit description of the Artin representation $\rho_{{\rm End}}\colon{\rm Gal}_{k}\to{\rm Aut}(\mathcal{O})\hookrightarrow{\rm GL}_{4}({\rm\mathbb{Q}})$ describing the Galois action on the endomorphism ring $\mathcal{O}$ (Corollary 6.6);
  4. (4)
    relate the moduli space of bielliptic Picard curves to a unitary Shimura curve $Y/{\rm\mathbb{Q}}$ and a quaternionic Shimura curve (Section 4.3);
  5. (5)
    determine the finitely many $\bar{\rm\mathbb{Q}}$-isomorphism classes of Pryms $P/{\rm\mathbb{Q}}$ which are geometrically non-simple (Proposition 6.18), corresponding to the CM points on $Y$.
Our Torelli result (1) is perhaps surprising because Barth [4] has shown that the Prym construction on all bielliptic genus $3$ curves does not satisfy a Torelli theorem, but has one-dimensional fibres. There are other instances of algebraic families of abelian surfaces with QM in the literature [3, 21], but (2) is interesting because of the very simple nature of this family. Moreover, contrary to previous approaches, our proof constructs the quaternionic action explicitly and geometrically, using the automorphisms and the polarization of $P$. Part (3) gives an interesting source of abelian surfaces with large endomorphism field and, in the notation of [17], produces the first published examples of geometrically simple abelian surfaces with Sato–Tate group $J(E_{3})$ and $J(E_{6})$ (Example 6.7). Part (4) and Shimura reciprocity allow us to calculate all geometrically non-simple Prym varieties in (5).
1.2. Rational torsion subgroups
To further illustrate the rich and accessible nature of these curves, we classify the finite torsion subgroups that arise in the Mordell-Weil groups $P({\rm\mathbb{Q}})$ of Pryms of bielliptic Picard curves over ${\rm\mathbb{Q}}$.
Theorem 1.1.
Let $P/{\rm\mathbb{Q}}$ be the Prym surface of a bielliptic Picard curve $C/{\rm\mathbb{Q}}$. Then
$P({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq\begin{cases}{\rm\mathbb{Z}}/n{\rm\mathbb{Z}}&\mbox{ for some }n\in\{1,2,3,6\},\mbox{ or }\\ {\rm\mathbb{Z}}/n{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/n{\rm\mathbb{Z}}&\mbox{ for some }n\in\{2,3\},\mbox{ or }\\ {\rm\mathbb{Z}}/6{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}.&\\ \end{cases}$
Conversely, for each finite abelian group $G$ above, there exist infinitely many $\bar{{\rm\mathbb{Q}}}$-isomorphism classes of bielliptic Picard Prym surfaces $P/{\rm\mathbb{Q}}$ with $P({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq G$.
Theorem 1.1 is the analogue of Mazur’s classification of rational torsion points of elliptic curves [33] in our setting. As with elliptic curves, each finite group allowed by Theorem 1.1 can be realized by an explicit family (in fact, sometimes multiple families) of Pryms whose coefficients can be rationally parameterized. This is in accordance with the philosophy of Mazur and Ogg, namely that torsion should only occur as a consequence of the ambient geometry of the relevant moduli space; see [34] for a recent survey. Theorem 1.1 is significantly easier to prove than Mazur’s theorem, essentially because $P$ has everywhere potentially good reduction. Nonetheless, it appears to be the first classification of rational torsion on a universal abelian variety over a Shimura variety which is not a modular curve.
Remark 1.2.
In a separate work with Schembri and Voight [28], we significantly extend some of our arguments here to show that for any abelian surface $A/{\rm\mathbb{Q}}$ such that ${\rm End}(A_{\bar{{\rm\mathbb{Q}}}})$ is a maximal order in a non-split quaternion algebra over ${\rm\mathbb{Q}}$, there is the uniform bound $|A({\rm\mathbb{Q}})_{\mathrm{tors}}|\leq 18$. Theorem 1.1 shows that this bound is sharp. Contrary to [28, Theorem 1.3], Theorem 1.1 is a complete classification, which we achieve using arguments specific to the geometry of the abelian surfaces considered here.
Example 1.3.
By Proposition 7.16 below, for each $t\in{\rm\mathbb{Q}}\setminus\{0,1,-1\}$ the Prym $P_{t}$ associated to the curve
$2(t^{2}-1)^{2}y^{3}=(x^{2}-1)(x^{2}-t^{2})$
satisfies $P_{t}({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$, achieving the maximum order allowed by Theorem 1.1.
Remark 1.4.
Theorem 1.1 implies that the rational torsion subgroup $J({\rm\mathbb{Q}})_{\mathrm{tors}}$ of the Jacobian of a bielliptic Picard curve is $12$-torsion. A complete classification of the groups $J({\rm\mathbb{Q}})_{\mathrm{tors}}$ would require parameterizing points of order $4$ in $J({\rm\mathbb{Q}})$.
It is natural to ask how the torsion subgroup of $P({\rm\mathbb{Q}})$ is affected by the presence of endomorphisms defined over ${\rm\mathbb{Q}}$. Generically, we have ${\rm End}(P)={\rm\mathbb{Z}}$, but it can also happen that ${\rm End}(P)$ is an order in a quadratic field. Such $P$ are said to be of ${\rm GL}_{2}$-type, and are modular by work of Khare-Wintenberger [25]. After proving explicit conditions on $a$ and $b$ for $P$ to be of ${\rm GL}_{2}$-type (see §6.3), we classify all the possible ${\rm End}(P)$-modules that arise as $P({\rm\mathbb{Q}})_{\mathrm{tors}}$.
Theorem 1.5.
Let $P/{\rm\mathbb{Q}}$ be the Prym surface of a non-CM bielliptic Picard curve over ${\rm\mathbb{Q}}$, with ${\rm End}(P)\neq{\rm\mathbb{Z}}$. Then ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{D}]$ for some $D\in\{2,6\}$. Furthermore, there is an isomorphism of ${\rm\mathbb{Z}}[\sqrt{D}]$-modules
$P({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq\begin{cases}\{0\}&\mbox{ or }\\ {\rm\mathbb{Z}}[\sqrt{D}]/\mathfrak{a_{p}}&\mbox{ for some $p\in\{2,3\}$ },\end{cases}$
where $\mathfrak{a}_{p}$ is the unique prime ideal of ${\rm\mathbb{Z}}[\sqrt{D}]$ above $p$. Conversely, for each $D\in\{2,6\}$, and for each of the three cyclic ${\rm\mathbb{Z}}[\sqrt{D}]$-modules $G$ that appear above, there are infinitely many $\overline{{\rm\mathbb{Q}}}$-isomorphism classes of Pryms $P/{\rm\mathbb{Q}}$ such that ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{D}]$ and $P({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq G$ as ${\rm\mathbb{Z}}[\sqrt{D}]$-modules.
When $D=2$ and $p=3$, Theorem 1.5 gives examples of ${\rm GL}_{2}$-type Pryms such that $P({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq{\rm\mathbb{F}}_{9}$, showing that the upper bound proven in [28, Theorem 1.4] for the order of the rational torsion subgroup of a QM abelian surface of ${\rm GL}_{2}$-type over ${\rm\mathbb{Q}}$ is in fact sharp.
1.3. Methods
The geometry of $(1,2)$-polarized surfaces was analyzed in detail by Barth [4] and his results are crucial to our study of $C$ and $P$. The key player in the story is the bigonal dual curve $\widehat{C}\colon y^{3}=x^{4}+8ax^{2}+16(a^{2}-4b)$, which is the intersection of $P$ with a well chosen theta divisor on $J$. The corresponding line bundle $\mathscr{M}=\mathcal{O}_{P}(\widehat{C})$ represents the $(1,2)$-polarization on $P$. The linear system $|\mathscr{M}|$ is a pencil of genus $3$ curves, and the $\mu_{3}$-action on $C$ induces a $\mu_{3}$-action on $P$ and on the pencil. The technical heart of this paper is to analyze this $\mu_{3}$-action on the pencil $|\mathscr{M}|$, both theoretically and explicitly. Our Torelli theorem (1) boils down to showing that $\widehat{C}$ is the unique bielliptic Picard curve in $|\mathscr{M}|$ of the correct signature. For the QM property (2), we give two different proofs. The first is indirect, relating the moduli of bielliptic Picard curves to a unitary Shimura curve (in other words, we first prove (4) and then deduce (2)). The second proof constructs the QM explicitly, by showing that a specific sextic twist of the original curve $C$ also lives in $|\mathscr{M}|$, but carrying the opposite signature. This second proof allows us to prove (3) as well.
To prove Theorems 1.1 and 1.5 we must first exhibit explicit families with specified torsion subgroups, and then we must rule out all other finite abelian groups. To exhibit groups, we describe elements of $P[2]$ and $P[3]$ geometrically, as coming from the fixed points for the $\mu_{6}$-action. To rule out groups which are not $6$-torsion, we first use general arguments for QM abelian surfaces which are developed in greater generality in [28]. Instead of quoting the main results there, we give a simpler proof (tailored to our special family of surfaces), which also allows us to handle the geometrically non-simple case not treated in [28]. [^1] For those interested in reading that paper as well, it is perhaps instructive to first read §8 of this paper (consulting [28] for the details of a few key results), before reading the more elaborate proofs in [28], where nothing is assumed about the geometry of the abelian surfaces involved. However, these general arguments only go so far; for example, they are not enough to rule out $4$-torsion in general QM abelian surfaces $A/{\rm\mathbb{Q}}$. To eliminate groups such as ${\rm\mathbb{Z}}/4{\rm\mathbb{Z}}$ and $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\times({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})$, we use arguments that are very much specific to the geometry of bielliptic Picard curves (see Propositions 8.5 and 8.6).
1.4. Previous results
Petkova–Shiga [36] considered bielliptic Picard curves from a complex analytic viewpoint. Using period matrix computations, they show that their Pryms over ${\rm\mathbb{C}}$ have quaternionic multiplication. Hashimoto-Murabayashi [21] and Baba-Granath [3] have studied genus two curves whose Jacobians $J^{\prime}$ have QM by $\mathcal{O}$. Over $\overline{k}$, each surface $J^{\prime}$ becomes isomorphic to a bielliptic Picard Prym $P$, but generally not over $k$. There is also work of Bonfanti-van Geemen [6] which gives an algebraic description of the embedding of the Shimura curve inside the moduli space of $(1,2)$-polarized abelian surfaces.
1.5. Future work
There are many other avenues of research related to bielliptic Picard curves and their Pryms that are ripe for investigation. In [29], we use our description of the ${\rm Gal}_{k}$-module ${\rm End}(P_{k^{\operatorname{sep}}})$ to characterize the bielliptic Picard curves with torsion Ceresa cycle. One can also try to study the average rank of the Mordell-Weil group $P({\rm\mathbb{Q}})$; the average rank of $P({\rm\mathbb{Q}})$ in cubic and sextic twist families was already considered in [1] and [42]. To help explore other arithmetic questions, it would be worthwhile to develop more robust algorithms that are tailored towards these curves, such as explicit descent algorithms, explicit addition laws, an analogue of Tate’s algorithm, $L$-functions, etc. We hope this work encourages others to study some of these topics.
1.6. Structure of paper
We start with some generalities on bielliptic Picard curves and their (dual) Prym varieties in §2. In particular, we discuss bigonal duality in §2.7. We prove our Torelli theorem in §3. In §4 we make the connection with unitary and quaternionic Shimura curves. In §5 we perform some explicit calculations with a pencil of curves on the Prym variety. These calculations are used in §6 to explicitly construct the quaternionic multiplication. In §7 we give explicit descriptions of the $2$ and $3$-torsion subgroups of Pryms of bielliptic Picard curves. In §8 and §9 we classify rational torsion subgroups and prove Theorems 1.1 and 1.5.
1.7. Acknowledgements
We thank Adam Morgan, Ciaran Schembri, Michael Stoll, and John Voight for helpful conversations and remarks. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157). Part of this research was carried out while the first author was visiting the second author in Jerusalem. We thank the Hebrew University of Jerusalem and the Einstein Institute of Mathematics for their hospitality.
1.8. Notation and conventions
  • Our base field will typically be denoted by $k$, with choice of separable and algebraic closure $k^{\operatorname{sep}}\subset\bar{k}$ and absolute Galois group ${\rm Gal}_{k}={\rm Gal}(k^{\operatorname{sep}}/k)$. All fields in this paper will be assumed to have characteristic $\neq 2,3$ unless explicitly stated otherwise. Galois actions will typically be right actions.
  • A variety over a field $k$ is a separated finite type scheme over $k$. A variety is called nice if it is smooth, projective and geometrically integral.
  • If $X,Y/k$ are varieties, $f\colon X_{k^{\operatorname{sep}}}\rightarrow Y_{k^{\operatorname{sep}}}$ is a morphism of $k^{\operatorname{sep}}$-varieties and $\sigma\in{\rm Gal}_{k}$, then $f^{\sigma}$ denotes the $k^{\operatorname{sep}}$-morphism $x\mapsto f(x^{\sigma^{-1}})^{\sigma}$.
  • For a nice variety $X/k$, denote by ${\rm Pic}(X)$ its Picard group and by ${\rm Pic}_{X}$ its Picard scheme.
  • If $X/k$ is a nice curve, let ${\rm Pic}^{n}(X)\subset{\rm Pic}(X)$ be the subset of line bundles of degree $n$, let $\operatorname{Jac}_{X}={\rm Pic}^{0}_{X}$ denote its Jacobian, let ${\rm Aut}(X)$ denote the $k$-automorphism group of $X$ and let $\mathbf{Aut}(X)$ be the scheme of automorphisms of $X$, with $\mathbf{Aut}(X)(K)={\rm Aut}(X_{K})$ for every field extension $K/k$.
  • For every integer $n\geq 1$ we define the group scheme $\mu_{n}={\rm Spec}\left(k[t]/(t^{n}-1)\right)$. If $X/k$ a variety, we define a $\mu_{n}$-action on $X$ to be a morphism of $k$-schemes $\mu_{n}\times_{k}X\rightarrow X$ satisfying the axioms of a (left) group action. If $m,n$ are coprime, giving a $\mu_{mn}$-action is the same as giving commuting $\mu_{m}$ and $\mu_{n}$-actions, using the isomorphism $\mu_{m}\times\mu_{n}\rightarrow\mu_{mn}$ induced by the inclusion maps.
  • We will typically write $\omega\in\bar{k}$ for a choice of primitive third root of unity and $k(\omega)$ for the smallest field extension of $k$ containing such $\omega$.
  • If $V$ is a vector space over a field $k$, we let ${\rm\mathbb{P}}(V)=\operatorname{Proj}({\rm Sym}^{\bullet}(V^{\vee}))$ be the projective space parametrizing lines in $V$. If $X/k$ is variety, a morphism $X\rightarrow{\rm\mathbb{P}}(V)$ is the same as a pair $(\mathscr{L},\phi)$, where $\mathscr{L}$ is a line bundle on $X$ and $\phi\colon V^{\vee}\otimes_{k}\mathcal{O}_{X}\rightarrow\mathscr{L}$ is a surjection (up to a suitable notion of isomorphism).
  • If $\mathscr{L}$ is a line bundle on a variety $X/k$ we denote by $|\mathscr{L}|:={\rm\mathbb{P}}(\operatorname{H}^{0}(X,\mathscr{L})^{\vee})$ the linear system of effective divisors in $\mathscr{L}$. If $D$ is a Cartier divisor on $X$ we also write $|D|$ for $|\mathcal{O}_{X}(D)|$. If $\mathscr{L}$ has no base points, we thus get a natural morphism $\phi_{\mathscr{L}}\colon X\rightarrow|\mathscr{L}|$.
  • If $A/k$ is an abelian variety, let $A^{\vee}={\rm Pic}^{0}_{A}$ be its dual abelian variety. If $\mathscr{L}\in{\rm Pic}(A)$, denote by $\lambda_{\mathscr{L}}\colon A\to A^{\vee}$ the homomorphism $x\mapsto t_{x}^{*}\mathscr{L}\otimes\mathscr{L}^{-1}$.
2. Bielliptic Picard curves and their Pryms
2.1. Basic definitions
Let $k$ be a field (always assumed of characteristic $\neq 2,3$).
Definition 2.1.
Let $C/k$ be a nice genus $3$ curve. We call a faithful $\mu_{6}$-action $\gamma\colon\mu_{6}\hookrightarrow\mathbf{Aut}(C)$ on $C$ bielliptic if the quotient $C/\mu_{2}$ is a genus one curve. We say $C$ is a bielliptic Picard curve over $k$ if it admits a bielliptic $\mu_{6}$-action.
For $a,b\in k$, consider the projective plane curve $C_{a,b}$ over $k$ with affine equation
(2.1) $C_{a,b}\colon y^{3}=x^{4}+ax^{2}+b.$
The curve $C_{a,b}$ has a unique ($k$-rational) point $\infty$ at infinity, and $C_{a,b}$ is smooth if and only if
$\Delta_{a,b}:=16b(a^{2}-4b)$
is nonzero. The $\mu_{2}$-action ±$\pm 1\cdot(x,y)=(\pm x,y)$ and $\mu_{3}$-action $\omega\cdot(x,y)=(x,\omega y)$ (where $\omega^{3}=1$) combine to a faithful bielliptic $\mu_{6}$-action $\gamma_{a,b}$, given by $\zeta\cdot(x,y)=(\zeta^{3}x,\zeta^{4}y)$, for every $\zeta\in\mu_{6}(k^{\operatorname{sep}})$. The unique fixed point for this action is $\infty$.
Theorem 2.2.
Let $C/k$ be nice genus $3$ curve with faithful $\mu_{6}$-action $\gamma$. Then the following conditions are equivalent:
  1.    $(1)$
    $\gamma$ is a bielliptic $\mu_{6}$-action (and consequently $C$ is a bielliptic Picard curve).
  2.    $(2)$
    $\gamma$ has a unique fixed point.
  3.    $(3)$
    $C/\mu_{3}$ has genus zero.
  4.    $(4)$
    $C$ is a plane curve.
  5.    $(5)$
    There is an isomorphism $C\simeq C_{a,b}$ for some $a,b\in k$ such that $\gamma$ corresponds to $\gamma_{a,b}$ or $\gamma_{a,b}^{-1}$.
Proof.
For each $i$ dividing $6$, let $n_{i}$ be the number of $\mu_{i}$-fixed points. By Riemann-Hurwitz, we have $n_{2}\in\{0,4,8\}$, $n_{3}\in\{2,5\}$, and $n_{6}\leq 3$. Moreover, $\mu_{2}$ permutes the $\mu_{3}$-fixed points and vice versa. So $n_{2}-n_{6}$ is divisible by $3$ and $n_{3}-n_{6}$ is divisible by $2$. There are then three possibilities:
  1.    $(a)$
    $(n_{2},n_{3},n_{6})=(0,2,0)$, or
  2.    $(b)$
    $(n_{2},n_{3},n_{6})=(4,5,1)$, or
  3.    $(c)$
    $(n_{2},n_{3},n_{6})=(8,2,2)$.
The equivalence of $(1)$, $(2)$, and $(3)$ follows immediately from the above cases. To show the equivalence of $(1)$ and $(4)$, first observe that case $(c)$ is hyperelliptic because $C/\mu_{2}$ is genus 00. Case $(a)$ is also hyperelliptic since an unramified double cover of a genus two curve is again hyperelliptic [16]. It remains to show that $C$ is non-hyperelliptic in case $(b)$. In that case, there is only one $\mu_{6}$-fixed point $\infty$, which is therefore $k$-rational. Moreover, $\infty$ must have different local monodromy type for the $\mu_{3}$-cover $C\to C/\mu_{3}\simeq{\rm\mathbb{P}}^{1}$ compared to the other four ramification points (which are permuted by the $\mu_{2}$-action). By Kummer theory, there is a model $C\colon y^{3}=f(x)$, for some quartic polynomial $f(x)\in k[x]$, and in particular $C$ is a plane quartic curve. This shows that $(1)-(4)$ are equivalent.
It is clear that $(5)$ implies all the rest, so it remains to show the converse. However, we have seen that such a $C$ has a model $y^{3}=f(x)$ which we may assume takes the form
(2.2) $y^{3}=x^{4}+ax^{2}+cx+d$
with the $\mu_{3}$-action $\omega\cdot(x,y)=(x,\omega y)$ or $\omega\cdot(x,y)=(x,\omega^{-1}y)$ for every $\omega\in\mu_{3}$. We see from this model (as in the proof of [10, Lemma 1.21(a)]) that $C$ admits an involution commuting with the $\mu_{3}$-action only if $c=0$ and this involution is given by $(x,y)\mapsto(-x,y)$, proving $(5)$. ∎
Remark 2.3.
A Picard curve is a plane quartic curve of the form (2.2). Such a curve is bielliptic (admitting a $\mu_{3}$-equivariant double cover to an elliptic curve) precisely when $c=0$, explaining our terminology.
There is one $k^{\operatorname{sep}}$-isomorphism class of bielliptic Picard curves that behaves so differently that it deserves its own name.
Definition 2.4.
A bielliptic Picard curve $C$ is special if it is $k^{\operatorname{sep}}$-isomorphic to $C_{0,1}:y^{3}=x^{4}+1$.
A bielliptic Picard curve $C$ is special precisely when ${\rm Aut}(C_{k^{\operatorname{sep}}})$ is larger than ${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$; in that case, it is a group of order $48$ [30, Theorem 3.1], with GAP label (48,33). We will see in Lemma 2.9 that $C_{a,b}$ is special if and only if $a=0$.
2.2. Marked bielliptic Picard curves
It is often useful to pin down a $\mu_{6}$-action on a bielliptic Picard curve with specified signature around the (unique, by Theorem 2.2) $\mu_{6}$-fixed point $\infty$.
Definition 2.5.
A marked bielliptic Picard curve is a bielliptic Picard curve $C$ with bielliptic $\mu_{6}$-action $\gamma$ such that $\mu_{6}$ acts on the tangent space $T_{\infty}C$ via the identity character $\mu_{6}\rightarrow\mu_{6}$.
An isomorphism of marked bielliptic Picard curves $(C,\gamma)\rightarrow(C^{\prime},\gamma^{\prime})$ is by definition an isomorphism of curves $C\rightarrow C^{\prime}$ that is equivariant with respect to the $\mu_{6}$-actions. A calculation using the generator $\frac{x}{y}\in(T_{\infty}C_{a,b})^{\vee}$ shows that the pair $(C_{a,b},\gamma_{a,b})$ defined in §2.1 is a marked bielliptic Picard curve.
Lemma 2.6.
Let $C/k$ be a bielliptic Picard curve with bielliptic $\mu_{6}$-action $\gamma$. Then one of $(C,\gamma^{\pm 1})$ is a marked bielliptic Picard curve, and if $C$ is not special this marked $\mu_{6}$-action is the unique one.
Proof.
Since $\gamma$ is faithful, it acts on $T_{\infty}C$ via the identity character or its inverse, so either $\gamma$ or $\gamma^{-1}$ acts via the identity. If $C$ is not special, $\mathbf{Aut}(C)\simeq\mu_{6}$ by [30, Theorem 3.1], proving the uniqueness in that case. ∎
2.3. Automorphisms
Lemma 2.7.
Every isomorphism $(C_{a,b},\gamma_{a,b})\rightarrow(C_{a^{\prime},b^{\prime}},\gamma_{a^{\prime},b^{\prime}})$ of marked bielliptic Picard curves over $k$ is of the form $(x,y)\mapsto(\lambda^{3}x,\lambda^{4}y)$ for some $\lambda\in k^{\times}$ such that $(a^{\prime},b^{\prime})=(\lambda^{6}a,\lambda^{12}b)$.
Proof.
Let $\phi\colon C_{a,b}\rightarrow C_{a^{\prime},b^{\prime}}$ be an isomorphism preserving the $\mu_{6}$-actions. Since $C_{a,b}$ and $C_{a^{\prime},b^{\prime}}$ are canonically embedded, $\phi$ is induced by a linear isomorphism of the ambient projective space ${\rm\mathbb{P}}^{2}_{k}$. This isomorphism preserves the point at infinity (being the unique $\mu_{6}$-fixed point) and the $\mu_{6}$-eigenspaces of $\operatorname{H}^{0}({\rm\mathbb{P}}^{2},\mathcal{O}(1))=\text{span}\{1,x,y\}$. A short calculation shows that $\phi$ must be of the form above. ∎
Lemma 2.8.
Every marked bielliptic Picard curve over $k$ is of the form $(C_{a,b},\gamma_{a,b})$ for some $a,b\in k$. Two marked bielliptic Picard curves $(C_{a,b},\gamma_{a,b})$ and $(C_{a^{\prime},b^{\prime}},\gamma_{a^{\prime},b^{\prime}})$ are isomorphic if and only if there exists $\lambda\in k^{\times}$ such that $a^{\prime}=\lambda^{6}a$ and $b^{\prime}=\lambda^{12}b$.
Proof.
The first part follows from Theorem 2.2(5). The second part follows from Lemma 2.7. ∎
Lemma 2.9.
The bielliptic Picard curve $C_{a,b}$ is special if and only if $a=0$.
Proof.
We may assume that $k$ is algebraically closed. Lemma 2.8 shows that $C_{a,b}$ is special if $a=0$. Conversely, suppose that $C_{a,b}$ is special. By [30, Theorem 3.1], $G={\rm Aut}(C_{a,b})$ is a group of order $48$ with GAP label (48,33). Moreover, ${\rm Aut}(C_{a,b},\gamma_{a,b})$ is the centralizer in $G$ of a cyclic order $6$ subgroup. A group theory calculation similar to [10, Lemma 4.1.5(c)] shows that all cyclic order $6$ subgroup of $G$ are conjugate and have centralizer of order $12$, so ${\rm Aut}(C_{a,b},\gamma_{a,b})$ has order $12$. On the other hand, Lemma 2.7 shows that ${\rm Aut}((C_{a,b},\gamma_{a,b}))=\{\lambda\in k^{\times}\mid(a,b)=(\lambda^{6}a,\lambda^{12}b)\}$. If $a\neq 0$, the latter set has size $6$, so if $C$ is special then $a=0$. ∎
Lemma 2.10.
Let $(C,\gamma)$ be a marked bielliptic Picard curve over $k$. Then $\mathbf{Aut}(C,\gamma)=\mu_{6}$ if $C$ is not special and $\mathbf{Aut}(C,\gamma)=\mu_{12}$ if $C$ is special.
Proof.
Combine Lemmas 2.7 and 2.9. ∎
Lemma 2.8 shows that the moduli stack of marked bielliptic Picard curves can be identified with the weighted projective stack ${\rm\mathbb{P}}(6,12)$ minus the discriminant locus $\Delta_{a,b}=0$. As with the moduli stack of elliptic curves, the coarse space has a simpler description. Given a bielliptic Picard curve $C/k$, its $j$-invariant is defined as
(2.3) $j(C):=j_{a,b}:=\frac{4b-a^{2}}{4b}\in k\setminus\{0\},$
where $a,b\in k$ are such that $C\simeq C_{a,b}$. Lemmas 2.6, 2.8 and 2.9 show that this is well defined and that $C_{k^{\operatorname{sep}}}\simeq C^{\prime}_{k^{\operatorname{sep}}}$ if and only if $j(C)=j(C^{\prime})$, even if $C$ or $C^{\prime}$ is special.
2.4. Twists
Let $(C,\gamma)$ be a marked bielliptic Picard curve over $k$. By the twisting principle, every cocycle $\xi\in\operatorname{H}^{1}(k,\mathbf{Aut}(C,\gamma))$ determines a bielliptic Picard curve $(C_{\xi},\gamma_{\xi})$ that is $k^{\operatorname{sep}}$-isomorphic to $(C,\gamma)$. Its $k$-isomorphism class is characterized by the existence of a $k^{\operatorname{sep}}$-isomorphism $\phi\colon(C_{\xi},\gamma_{\xi})_{k^{\operatorname{sep}}}\rightarrow(C,\gamma)_{k^{\operatorname{sep}}}$ such that $\sigma\mapsto\phi^{\sigma}\circ\phi^{-1}$ represents the class of $\xi$.
In particular, since $\gamma$ determines a subgroup $\mu_{6}\subset\mathbf{Aut}(C,\gamma)$ and since $\operatorname{H}^{1}(k,\mu_{6})\simeq k^{\times}/k^{\times 6}$, every $\delta\in k^{\times}$ gives rise to a sextic twist $(C_{\delta},\gamma_{\delta})$ of $(C,\gamma)$. Concretely, if $C=C_{a,b}$ then the sextic twist of $C_{a,b}$ by $\delta$ is isomorphic to $C_{\delta a,\delta^{2}b}$. If $C$ is not special, Lemmas 2.6 and 2.7 show that $\mathbf{Aut}(C,\gamma)=\mathbf{Aut}(C)=\mu_{6}$ and so every twist of $C$ (as a marked and unmarked curve) is isomorphic to a sextic twist.
2.5. The Prym variety $P$
Let $(C,\gamma)$ be a marked bielliptic Picard curve over $k$. We associate to $C$ an abelian surface $P$ whose study will occupy much of the rest of this paper.
Restricting the $\mu_{6}$-action to $\mu_{2}$ defines an involution $\tau\colon C\rightarrow C$. The quotient $\pi\colon C\rightarrow E:=C/\tau$ is a genus $1$ curve. When endowed with the $k$-point $\pi(\infty)$ it has the structure of an elliptic curve, so we can identify $\operatorname{Jac}_{E}$ with $E$. Let $J=\operatorname{Jac}_{C}$ be the Jacobian of $C$.
Definition 2.11.
The Prym variety of $C$ is defined as $P:=\ker(1+\tau^{*}\colon J\rightarrow J)$.
This is an abelian surface. We refer to [27, §3.3] for more details concerning Prym varieties of bielliptic genus $3$ curves. The properties we will use are summarized in the following commutative diagram:
[…]
It depicts two dual exact sequences, where $\pi_{*}$ (resp. $\pi^{*}$) denotes pushforward (resp. pullback) of divisors. The principal polarization on $J$ restricts to a $(1,2)$-polarization $\lambda\colon P\rightarrow A:=P^{\vee}$ with kernel $P[\lambda](k^{\operatorname{sep}})\simeq({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}$. The intersection of $P$ and $\pi^{*}(E)$ is $\pi^{*}(E[2])=P[\lambda]$. The sum of the two maps $P\rightarrow J$ and $E\rightarrow J$ is an isogeny $P\times E\rightarrow J$ with kernel $\{(\pi^{*}(x),x)\mid x\in E[2]\}\simeq E[2]$.
If $C=C_{a,b}$ for some $a,b\in k$, we may write $E_{a,b}$, $J_{a,b}$, $\lambda_{a,b}\colon P_{a,b}\rightarrow A_{a,b}$ et cetera, but we drop the subscripts when convenient.
Remark 2.12.
$P$ need not be principally polarizable over $k$; indeed we prove in Corollary 6.14 that if $P$ is geometrically simple, then it is principally polarizable if and only if $16b(a^{2}-4b)$ is a sixth power in $k^{\times}$. However, if $b=s^{3}$ is a cube in $k^{\times}$, then the Jacobian of the following genus two curve is in the isogeny class of $P_{a,s^{3}}$:
$-asy^{2}=(x^{2}+2x-2)(s^{3}x^{4}+4s^{3}x^{3}+2dx-d).$
where $d=a^{2}-4s^{3}$; this follows from [39, Theorem 1.1].
We now incorporate the $\mu_{3}$-action on $C$ into the picture. By Albanese functoriality, there is a unique $\mu_{6}$-action on $J$ such that the Abel–Jacobi map $\text{AJ}_{\infty}\colon C\hookrightarrow J$ is $\mu_{6}$-equivariant. This induces a $\mu_{3}$-action on the subvariety $P$ and the quotient $E$. We record the following important signature calculation.
Lemma 2.13.
For every primitive third root of unity $\omega\in k^{\operatorname{sep}}$, the action of $\omega$ on $\operatorname{H}^{0}(P_{k^{\operatorname{sep}}},\Omega^{1}_{P_{k^{\operatorname{sep}}}})$ has characteristic polynomial $T^{2}+T+1$.
Proof.
We may assume that $(C,\gamma)=(C_{a,b},\gamma_{a,b})$ by Lemma 2.8 and that $k=k^{\operatorname{sep}}$, so fix such a $\omega\in k$. Then $\omega$ acts on $C_{a,b}$ via $(x,y)\mapsto(x,\omega y)$. The vector space $\operatorname{H}^{0}(C,\Omega^{1}_{C})$ has basis $\frac{dx}{y^{2}},x\frac{dx}{y^{2}},\frac{dx}{y}$, hence the $\omega$-action on this vector space has eigenvalues $\omega,\omega,\omega^{2}$. The vector space $\operatorname{H}^{0}(E,\Omega^{1}_{E})$ has basis $\frac{dx}{y^{2}}$, hence $\omega$ has eigenvalue $\omega$. Since the map $P\times E\rightarrow J,(p,e)\mapsto p+\pi^{*}(e)$ is an isogeny with kernel $E[2]$, it induces a $\mu_{3}$-equivariant isomorphism $\operatorname{H}^{0}(J,\Omega^{1}_{J})\simeq\operatorname{H}^{0}(P,\Omega^{1}_{P})\oplus\operatorname{H}^{0}(E,\Omega^{1}_{E})$. Combining the last three sentences with the isomorphism $\operatorname{H}^{0}(C,\Omega^{1}_{C})\simeq\operatorname{H}^{0}(J,\Omega_{J}^{1})$ proves the lemma. ∎
Lemma 2.14.
The subgroup of $\mu_{3}$-fixed points of $P$ is of size $9$ and contained in $P[3]$.
Proof.
We may assume that $k=k^{\operatorname{sep}}$. So let $\omega\in k$ be a third root of unity and $\alpha\colon P\rightarrow P$ the corresponding automorphism. Let $\beta=1-\alpha$. By Lemma 2.13, $\beta$ induces an isomorphism on differentials so $\beta$ is an isogeny. Since $\beta\circ(\alpha^{2}+\alpha+1)=0$ and $\beta$ is surjective,
(2.4) $\alpha^{2}+\alpha+1=0.$
Therefore $\beta^{2}=1-2\alpha+\alpha^{2}=-3\alpha$, so $\beta$ has degree $9$ and $P[\beta]\subset P[-3\alpha]=P[3]$. ∎
2.6. The dual Prym variety $A$
Keep the notations from §2.5. Since $P[\lambda]\subset P[2]$, there exists a unique isogeny [^2] We warn the reader that $\widehat{\lambda}$ is not the same as the dual $\lambda^{\vee}$ of $\lambda$. Being a polarization, $\lambda$ is in fact self-dual! $\widehat{\lambda}\colon A\rightarrow P$ such that $\widehat{\lambda}\circ\lambda=[2]$. The next proposition describes this isogeny geometrically and characterizes $A$ in terms of the curve $C$ with its involution $\tau\colon C\rightarrow C$.
Proposition 2.15.
Let $i\colon C\rightarrow A$ be the composite of the Abel–Jacobi map $C\rightarrow J$ with respect to $\infty\in C(k)$ and the projection map $J\rightarrow A$.
  1. (1)
    The morphism $i\colon C\hookrightarrow A$ is a closed embedding.
  2. (2)
    The divisor $i(C)$ is ample and induces the polarization $\widehat{\lambda}$.
  3. (3)
    If $B/k$ is an abelian surface and $j\colon C\hookrightarrow B$ a closed embedding mapping $\infty$ to 00 and such that $[-1]\circ j=j\circ\tau$, then there exists a unique isomorphism of abelian surfaces $\phi\colon A\rightarrow B$ such that $\phi\circ i=j$.
Proof.
This is due to Barth [4]; see [27, Proposition 3.5] for a detailed proof. ∎
Let ${\rm Sym}^{2}C$ be the symmetric square of $C$, a nice surface over $k$ parameterizing effective divisors of degree $2$ on $C$. In analogy with Jacobians of genus $2$ curves, we will describe the fibres of the surjective map $i^{(2)}\colon{\rm Sym}^{2}C\rightarrow A$ given by $p+p^{\prime}\mapsto i(p)+i(p^{\prime})$; this is due to Ikeda [23].
To this end, we define two involutions on ${\rm Sym}^{2}C$. The first is $\tau^{(2)}(p+p^{\prime})=\tau(p)+\tau(p^{\prime})$. The second involution $\kappa$ sends $p+p^{\prime}$ to the unique effective degree $2$ divisor linearly equivalent to $4\infty-p-p^{\prime}$; this is well defined by Riemann–Roch and the fact that $C$ is not hyperelliptic. Since $\tau(\infty)=\infty$, the involutions $\tau^{(2)}$ and $\kappa$ on ${\rm Sym}^{2}C$ commute. Finally, the double cover $\pi\colon C\rightarrow E$ induces a map $\pi^{*}\colon E\rightarrow{\rm Sym}^{2}C$, giving an embedding $E\simeq\pi^{*}(E)\hookrightarrow{\rm Sym}^{2}C$.
Proposition 2.16 (Ikeda).
The map $i^{(2)}\colon{\rm Sym}^{2}C\rightarrow A$ contracts $\pi^{*}(E)$ to the origin. Two points $D,D^{\prime}\in{\rm Sym}^{2}C$ not in $\pi^{*}(E)$ map to the same point of $A$ if and only if $D^{\prime}=\kappa(\tau^{(2)}(D))$, i.e. $D^{\prime}+\tau(D)\sim 4\infty$.
Proof.
This follows from the proof of [23, Lemma 3.1]. ∎
We see that $A$ is obtained from ${\rm Sym}^{2}C/\langle\tau^{(2)}\circ\kappa\rangle$ by contracting the rational curve $\pi^{*}(E)/\langle-1\rangle$. This description will be useful in Section 7.5, where we relate the bitangents of $C$ to the group $A[2]$.
2.7. Bigonal duality
Given a marked bielliptic Picard curve $(C,\gamma)$, it turns out that we can define another marked bielliptic Picard curve $(\widehat{C},\hat{\gamma})$ such that the role of the Prym variety and its dual are reversed: the Prym variety of $\widehat{C}$ is isomorphic to $A=P^{\vee}$. This is called the bigonal dual of $C$, originally defined by Pantazis [35] (inspired by ideas of Donagi [12]) and analyzed by Barth [4] in this specific situation; see [27, §3.4] for more details. Let $\Theta_{2\infty}\subset J$ be the image of the Abel–Jacobi map ${\rm Sym}^{2}C\rightarrow J$ sending $P+P^{\prime}$ to $P+P^{\prime}-2\infty$.
Definition 2.17.
The bigonal dual of $C$ is defined by $\widehat{C}:=P\cap\Theta_{2\infty}$.
The $\mu_{3}$-action on $P$ restricts to a $\mu_{3}$-action on $\widehat{C}$. Inversion $[-1]$ on $P$ restricts to an involution $\hat{\tau}$ on $\widehat{C}$. Using the isomorphism $\mu_{2}\times\mu_{3}\rightarrow\mu_{6}$ induced by the two inclusions, we obtain a $\mu_{6}$-action $\hat{\gamma}$ on $\widehat{C}$.
Lemma 2.18.
The pair $(\widehat{C},\hat{\gamma})$ is a marked bielliptic Picard curve.
Proof.
By [27, Lemma 3.5], $\widehat{C}$ is a nice genus $3$ curve and $\widehat{C}/\mu_{2}$ is a genus $1$ curve. Since the $\mu_{3}$-action on $P$ has only isolated fixed points (Lemma 2.14), the $\mu_{3}$-action on $\widehat{C}$ is nontrivial. Therefore $\hat{\gamma}$ is a faithful bielliptic action and $\widehat{C}$ is a bielliptic Picard curve. The origin $0\in P(k)$ is a $\mu_{6}$-fixed point of $\widehat{C}$, which is unique by Theorem 2.2.
It remains to show that $(\widehat{C},\widehat{\gamma})$ is a marked bielliptic Picard curve, in other words that $\mu_{6}$ acts on $T_{0}\widehat{C}$ via the identity character. Equivalently, we will show that $\mu_{6}$ acts on $(T_{0}\widehat{C})^{\vee}$ via the inverse of the identity character. Choose a uniformizer $t\in\mathcal{O}_{C,\infty}$. This induces a $k$-basis $\{t_{1},t_{2}\}$ of $(T_{(\infty,\infty)}(C\times C))^{\vee}=(T_{\infty}C)^{\vee}\oplus(T_{\infty}C)^{\vee}$ and an isomorphism of completed local rings $\widehat{\mathcal{O}}_{C\times C,(\infty,\infty)}\simeq k[[t_{1},t_{2}]]$. In turn, this induces an isomorphism $\widehat{\mathcal{O}}_{{\rm Sym}^{2}C,2\infty}\simeq k[[u,v]]$, where $u=t_{1}+t_{2}$ and $v=t_{1}t_{2}$. It follows that $(T_{2\infty}{\rm Sym}^{2}C)^{\vee}\simeq(T_{0}\Theta_{2\infty})^{\vee}$ has basis $\{u,v\}$ and $[-1]$ sends $u,v$ to $-u,v$ respectively. Therefore $T_{0}\widehat{C}=T_{0}\Theta_{2\infty}\cap T_{0}P$ has basis $\{u\}$ which indeed has the correct character. ∎
It follows that the objects defined in §2.5 also apply to $(\widehat{C},\hat{\gamma})$, giving $\widehat{J}=\operatorname{Jac}_{\widehat{C}}$, $\widehat{P}$, $\widehat{A}$, $\hat{\pi}\colon\widehat{C}\rightarrow\widehat{E}$ in this context. The inclusion $\widehat{C}\hookrightarrow P$ induces a homomorphism $\widehat{J}\rightarrow P$.
Proposition 2.19.
The homomorphism $\widehat{J}\rightarrow P$ factors through the projection $\widehat{J}\rightarrow\widehat{A}$ and induces an isomorphism $\widehat{A}\xrightarrow{\sim}P$ of $(1,2)$-polarized surfaces.
Proof.
See [27, Proposition 3.6]. The fact that the isomorphism preserves the polarizations means that the maps $A\rightarrow P$ and $\widehat{P}\rightarrow\widehat{A}$ that are both denoted by $\widehat{\lambda}$ are the same under these identifications. ∎
Choosing an isomorphism $(C,\gamma)\simeq(C_{a,b},\gamma_{a,b})$ for some $a,b\in k$ using Lemma 2.8, we can write down equations for $\widehat{C}$ [27, §3.4 Eq (3.7), (3.8)]:
(2.5) $C_{a,b}$ $:y^{3}=x^{4}+ax^{2}+b,$
(2.6) $E_{a,b}$ $:y^{2}=x^{3}+16(a^{2}-4b),$
(2.7) $\widehat{C}_{a,b}$ $:y^{3}=x^{4}+8ax^{2}+16(a^{2}-4b),$
(2.8) $\widehat{E}_{a,b}$ $:y^{2}=x^{3}+b.$
The coordinates on $E_{a,b}$ and $\widehat{E}_{a,b}$ are chosen so that the double covers $C_{a,b}\rightarrow E_{a,b}$ and $\widehat{C}_{a,b}\rightarrow\widehat{E}_{a,b}$ are given by $(x,y)\mapsto(4y,8x^{2}+4a)$ and $(x,y)\mapsto(y/4,x^{2}/8+a/2)$ respectively.
The third equation (2.7) shows that bigonal duality takes the following elegant form on the $j$-invariant (2.3):
Lemma 2.20.
Let $C$ be a bielliptic Picard curve over $k$. Then $j(\widehat{C})=1/j(C)$.
The quartic polynomial $\hat{f}:=x^{4}+8ax^{2}+16(a^{2}-4b)$ defining the bigonal dual curve can also be constructed from the Galois theory of the original quartic $f:=x^{4}+ax^{2}+b$. A calculation shows:
Lemma 2.21.
Suppose that $f$ has roots {±$\{\pm\alpha,\pm\beta\}$ in $k^{\operatorname{sep}}$. Then $\hat{f}$ has roots {±$\{\pm 2\alpha\pm 2\beta\}$.
This description of $\hat{f}$ will be useful when analyzing $3$-torsion in $P$ in §7.2.
3. A Torelli theorem
A bielliptic Picard curve $C$ gives rise to two $(1,2)$-polarized abelian surfaces with $\mu_{3}$-actions: the Prym surface $P$ and its dual $A$. We show that one can recover $C$ from $A$, together with its polarization and $\mu_{3}$-action; in other words, we prove a Torelli-type theorem (Theorem 3.13). We do this by studying the pencil of genus three curves in the linear system $|C|$ on $A$. In fact, we will prove that any indecomposable $(1,2)$-polarized abelian surface with compatible $\mu_{3}$-action is the dual Prym variety of a bielliptic Picard curve, so we work in this generality below.
3.1. Generalities on $(1,2)$-polarized surfaces
We recall and complement some results of Barth [4]. Barth works over ${\rm\mathbb{C}}$, but the proofs of the basic results quoted here are valid in arbitrary characteristic $\neq 2$.
Let $(A,\lambda)$ be a $(1,2)$-polarized abelian surface over $k$. We assume that $(A,\lambda)$ is indecomposable, that is, not isomorphic to a product of two elliptic curves $(E,O)\times(E^{\prime},2O^{\prime})$ as polarized abelian varieties. Recall our notations concerning linear systems and abelian varieties in §1.8.
Lemma 3.1.
Let $\mathscr{L}\in{\rm Pic}(A_{k^{\operatorname{sep}}})$ be a line bundle representing the polarization: $\lambda_{\mathscr{L}}=\lambda$.
  1. (1)
    $\dim_{k}\operatorname{H}^{0}(A,\mathscr{L})=2$.
  2. (2)
    The linear system $|\mathscr{L}|$ has four base-points on which $A[\lambda](k^{\operatorname{sep}})$ acts simply transitively.
  3. (3)
    At most finitely many members of the pencil are singular (arithmetic) genus $3$ curves, in which case they are curves of geometric genus $2$ with one node or the union of two genus $1$ curves $X_{1},X_{2}$ with intersection number $(X_{1}\cdot X_{2})=2$.
Proof.
See [4, §1.2]. ∎
Lemma 3.2.
There exists a unique $\mathscr{L}\in{\rm Pic}(A_{k^{\operatorname{sep}}})$ representing $\lambda$ whose base locus is $A[\lambda]$. This line bundle is symmetric: $[-1]^{*}\mathscr{L}\simeq\mathscr{L}$. Moreover, for every divisor $D\in|\mathscr{L}|$ we have $[-1]^{*}D=D$.
Proof.
Denote the base locus of a line bundle $\mathscr{L}$ by $\text{BL}(\mathscr{L})$. Given $a\in A(k^{\operatorname{sep}})$, let $t_{a}\colon A\rightarrow A$ be the translation map and $\lambda(a)\in A^{\vee}(k^{\operatorname{sep}})={\rm Pic}^{0}(A_{k^{\operatorname{sep}}})$ be its image under $\lambda$. Then for any $\mathscr{L}$ representing $\lambda$, the following identity holds:
(3.1) $\text{BL}(\lambda(a)\otimes\mathscr{L})=t_{a}^{*}\text{BL}(\mathscr{L}).$
Indeed, by definition $\lambda(a)\otimes\mathscr{L}=t_{a}^{*}\mathscr{L}$. Let $D,D^{\prime}$ be two distinct members of the pencil $|\mathscr{L}|$. Then $t_{a}^{*}D,t_{a}^{*}D^{\prime}$ are two distinct members of $|t_{a}^{*}\mathscr{L}|$. So $\text{BL}(t_{a}^{*}\mathscr{L})=t_{a}^{*}D\cap t_{a}^{*}D^{\prime}=t_{a}^{*}(D\cap D^{\prime})=t_{a}^{*}\text{BL}(\mathscr{L})$, proving (3.1). This identity, together with Lemma 3.1(2), proves the existence and uniqueness of $\mathscr{L}$. The remainder of the lemma follows from [4, Proposition 1.6]. ∎
For the remainder of this section fix $\mathscr{L}\in{\rm Pic}(A_{k^{\operatorname{sep}}})$ satisfying the properties of Lemma 3.2. Since there is a unique such line bundle up to isomorphism and $A$ has a $k$-point, $\mathscr{L}$ is automatically Galois invariant and descends to a unique line bundle on $A$ that we will continue to denote by $\mathscr{L}$.
Remark 3.3.
This marks an interesting contrast with principal polarizations on abelian surfaces, which may not be represented by a $k$-rational line bundle. Indeed, this is the case for a positive proportion of genus 2 Jacobians over ${\rm\mathbb{Q}}$ ordered by height [37, Theorem 23].
The next proposition shows that every smooth member of the pencil $|\mathscr{L}|$ realizes $A$ as the dual Prym variety of a bielliptic genus $3$ curve, as in Proposition 2.15. A bielliptic genus $3$ curve $(X,\tau)$ is a nice genus $3$ curve together with an involution $\tau$ on $X$ with $4$ fixed points. For such a curve, let $\operatorname{Prym}(X,\tau)=\ker(1+\tau^{*}\colon\operatorname{Jac}_{X}\rightarrow\operatorname{Jac}_{X})$ be its Prym variety.
Proposition 3.4.
Suppose that $X\in|\mathscr{L}|(k)$ is smooth, and let $\tau$ be the restriction of $[-1]$ to $X$. Then $(X,\tau)$ is a bielliptic genus $3$ curve and the inclusion $X\rightarrow A$ induces an isomorphism $\operatorname{Prym}(X,\tau)^{\vee}\simeq A$ of $(1,2)$-polarized surfaces.
Proof.
See [4, Proposition (1.8) and Theorem (1.12)]. ∎
The abelian surface $A$ can also be reconstructed from the singular members of the pencil. By Lemma 3.1(3), there are two cases to consider.
For the first case, let $X\in|\mathscr{L}|(k)$ be a curve with one node, whose normalization $\tilde{X}\rightarrow X$ is a genus $2$ curve. Let $p\in X(k)$ be the node and $q_{1},q_{2}$ be its preimages in $\tilde{X}(k^{\operatorname{sep}})$. The morphism $\tilde{X}\rightarrow X\hookrightarrow A$ induces a homomorphism of abelian varieties $\phi\colon\operatorname{Jac}_{\tilde{X}}\rightarrow A$, sending a degree zero divisor class on $\tilde{X}$ to its sum in $A$.
Proposition 3.5.
The homomorphism $\phi\colon\operatorname{Jac}_{\tilde{X}}\rightarrow A$ is an isogeny with kernel $\{0,q_{1}-q_{2}\}$.
Proof.
The image of $\phi$ is an abelian subvariety containing $X$, so $\phi$ is surjective. By construction, $\phi(q_{1}-q_{2})=0$. It suffices to prove $\phi$ has degree $2$. This follows from an intersection number calculation. Given divisors $D,E$ on a nice surface, denote by $[D],[E]$ their class in the Picard group and their intersection number by $(D\cdot E)$. The origin $0\in A(k)$ lifts to a unique point $\infty\in\tilde{X}(k)$ and we identify $\tilde{X}$ with its image in $\operatorname{Jac}_{\tilde{X}}$ using the point $\infty$. Then $\phi^{*}([X])=(\deg\phi)[\tilde{X}]$ and by the projection formula
$(\deg\phi)(X,X)=(\phi^{*}X,\phi^{*}X)=(\deg\phi)^{2}(\tilde{X},\tilde{X}).$
The adjunction formula tells us that $(X,X)=2\cdot 3-2=4$ and $(\tilde{X},\tilde{X})=2$. Therefore $\deg\phi=2$, as claimed. ∎
For the second case, let $X=X_{1}\cup X_{2}\in|\mathscr{L}|(k)$ be a pair of genus $1$ curves with intersection number $2$ and let $\tilde{X}=X_{1}\sqcup X_{2}\rightarrow X$ be its normalization. The map $\tilde{X}\rightarrow X\hookrightarrow A$ induces a homomorphism $\phi\colon\operatorname{Jac}_{\tilde{X}}:=\operatorname{Jac}_{X_{1}}\times\operatorname{Jac}_{X_{2}}\rightarrow A$.
Proposition 3.6.
The map $\phi$ is an isogeny of degree $2$.
Proof.
The proof is very similar to that of Proposition 3.5; we omit the details. ∎
3.2. Compatible $\mu_{3}$-actions on $(1,2)$-polarized abelian surfaces
Definition 3.7.
A $\mu_{3}$-abelian surface over $k$ is a triple $(A,\lambda,\alpha)$, where $(A,\lambda)$ is a $(1,2)$-polarized abelian surface and $\alpha\colon\mu_{3}\rightarrow{\rm Aut}(A_{k^{\operatorname{sep}}})$ is a ${\rm Gal}_{k}$-equivariant group homomorphism such that for every primitive third root of unity $\omega\in k^{\operatorname{sep}}$:
  1. (1)
    $\alpha(\omega)^{\vee}\circ\lambda\circ\alpha(\omega)=\lambda$;
  2. (2)
    $\alpha(\omega)$ has characteristic polynomial $X^{2}+X+1$ on $T_{0}A=\operatorname{H}^{0}(A,\Omega^{1}_{A})^{\vee}$.
The signature condition (2) is equivalent to $\alpha(\omega)^{2}+\alpha(\omega)+1=0$, and is motivated by Lemma 2.13. Condition (1) can be reformulated as follows: after extending $\mu_{3}\hookrightarrow{\rm Aut}(A_{k^{\operatorname{sep}}})$ to a ${\rm\mathbb{Q}}$-algebra homomorphism ${\rm\mathbb{Q}}(\omega)\hookrightarrow{\rm End}(A_{k^{\operatorname{sep}}})\otimes{\rm\mathbb{Q}}$, the Rosati involution associated to $\lambda$ restricts to the conjugation involution $a+b\omega\mapsto a+b\bar{\omega}$ on ${\rm\mathbb{Q}}(\omega)$.
Remark 3.8.
It would be more accurate to call such triples ‘$(1,2)$-polarized abelian surfaces with compatible $\mu_{3}$-action of signature $(1,1)$’, but we stick with $\mu_{3}$-abelian surfaces for brevity.
Let $(A,\lambda,\alpha)$ be a $\mu_{3}$-abelian surface over $k$. Assume that $(A,\lambda)$ is indecomposable, so the results of §3.1 apply. Let $\mathscr{L}$ be the unique line bundle on $A$ representing $\lambda$ with base locus $A[\lambda]$. The compatibility of $\alpha$ and $\lambda$ and the uniqueness of $\mathscr{L}$ show that $\alpha(\omega)^{*}\mathscr{L}\simeq\mathscr{L}$ for all $\omega\in\mu_{3}(k^{\operatorname{sep}})$. Each such isomorphism is unique up to $k^{\times}$-scaling so determines a canonical isomorphism $|\alpha(\omega)^{*}\mathscr{L}|\simeq|\mathscr{L}|$ Via pullback this induces a $\mu_{3}$-action on the pencil $|\mathscr{L}|$.
Proposition 3.9.
Every $\mu_{3}$-fixed point of $|\mathscr{L}|$ is a smooth curve.
Proof.
We may assume that $k=k^{\operatorname{sep}}$. Let $\omega\in k$ be a third root of unity and for ease of notation denote $\alpha(\omega)\colon A\rightarrow A$ by $\alpha$. Let $X\in|\mathscr{L}|$ be fixed by $\alpha$. Assume for the sake of contradiction that $X$ is singular, with normalization $\tilde{X}\rightarrow X$. By Lemma 3.1(3) and Propositions 3.5 and 3.6, the map $\tilde{X}\rightarrow A$ induces an isogeny $\phi\colon\operatorname{Jac}_{\tilde{X}}\rightarrow A$ whose kernel $\{0,D\}$ has order $2$. Since $\alpha$ preserves $X$, it induces an automorphism $\tilde{\alpha}$ of $\operatorname{Jac}_{\tilde{X}}$ and $\phi$ is equivariant with respect to $\tilde{\alpha}$ and $\alpha$. It follows that $\tilde{\alpha}$ fixes $D$. Since $\alpha^{2}+\alpha+1=0$ (Condition (2) in Definition 3.7) and $\phi$ is an isogeny, $\tilde{\alpha}^{2}+\tilde{\alpha}+1=0$. Therefore the degree of $1-\tilde{\alpha}$ is a power of $3$, contradicting the fact that ${\rm\mathbb{Z}}/2{\rm\mathbb{Z}}\simeq\{0,D\}\subset\operatorname{Jac}_{\tilde{X}}[1-\tilde{\alpha}]$. ∎
Corollary 3.10.
Assume that $k=k^{\operatorname{sep}}$. Then there are exactly $2$ curves in $|\mathscr{L}|$ that are preserved by the $\mu_{3}$-action.
Proof.
Let $\omega\in k$ be a third root of unity. By Proposition 3.9, every fixed point of $\alpha(\omega)$ on $|\mathscr{L}|$ is smooth. Since $|\mathscr{L}|$ always contains singular curves [4, §1.2], this implies that $\alpha(\omega)$ acts nontrivially on $|\mathscr{L}|$. By the classification of order $3$ elements of ${\rm PGL}_{2}(k)$, we may choose coordinates on ${\rm\mathbb{P}}^{1}$ so that $\alpha(\omega)$ is given by $(x:y)\mapsto(\omega x:y)$. This map clearly has $2$ fixed points. ∎
Let $X_{1},X_{2}\in|\mathscr{L}_{k^{\operatorname{sep}}}|$ be the members of the pencil that are preserved by the $\mu_{3}$-action $\alpha$. By Lemma 3.2, the involution $[-1]$ also preserves $X_{1}$ and $X_{2}$. These combine to a $\mu_{6}$-action $\gamma_{i}$ on $X_{i}$ for each $i=1,2$. The next proposition shows that each $X_{i}$ is a bielliptic Picard curve and the $\mu_{6}$-actions have ‘opposite’ signature.
Proposition 3.11.
Reordering $X_{1}$ and $X_{2}$ if necessary, the pairs $(X_{1},\gamma_{1})$ and $(X_{2},\gamma_{2}^{-1})$ are marked bielliptic Picard curves. Moreover, these pairs are defined over $k$.
Proof.
The second sentence follows from the first, since ${\rm Gal}_{k}$ preserves the $\mu_{3}$-fixed points of $|\mathscr{L}_{k^{\operatorname{sep}}}|$ and the signatures of the $\mu_{6}$-action. To show that $(X_{1},\gamma_{1})$ and $(X_{2},\gamma_{2})$ are bielliptic Picard curves of opposite signature, we may assume that $k=k^{\operatorname{sep}}$.
Let $(X,\gamma)=(X_{i},\gamma_{i})$ for some $i=1,2$. Proposition 3.9 shows that $X$ is smooth, and Proposition 3.4 shows that $(X,[-1]|_{X})$ is a bielliptic genus 3 curve. Since $\alpha^{2}+\alpha+1=0$, the $\mu_{3}$-action on $A$ has isolated fixed points, so the induced $\mu_{3}$-action on $X$ is faithful. It follows that $\gamma$ is a faithful bielliptic action and hence $X$ is a bielliptic Picard curve. Note that $0\in X(k)\subset A(k)$ is the unique $\mu_{6}$-fixed point.
To prove the claim about signatures, it suffices to show that the $\mu_{3}$-action on $T_{0}X_{1}$ and $T_{0}X_{2}$ act via inverse characters. Since $\mathscr{L}$ has self-intersection $4$ and $|\mathscr{L}|$ has four base points (Lemma 3.1), $X_{1}$ and $X_{2}$ intersect transversally at all these base points. Since $0\in A(k)$ is such a base point,
(3.2) $T_{0}A=T_{0}X_{1}\oplus T_{0}X_{2}.$
Condition (2) of Definition 3.7 says that the $\mu_{3}$-action on $T_{0}A$ is a direct sum of the two nontrivial characters of $\mu_{3}$. Since the decomposition (3.2) is $\mu_{3}$-stable, the proposition follows. ∎
3.3. A Torelli theorem for marked bielliptic Picard curves
Let $(C,\gamma)$ be a marked bielliptic Picard curve over $k$. The $\mu_{6}$-action on $C$ induces a $\mu_{6}$-action on $J=\operatorname{Jac}_{C}$ such that the Abel–Jacobi map $C\hookrightarrow J$ is $\mu_{6}$-equivariant. This induces $\mu_{3}$-actions $\alpha$ on $P$ and $\hat{\alpha}$ on [^3] A calculation shows that the $\mu_{3}$-actions $\hat{\alpha}$ (making $J\rightarrow A$ $\mu_{3}$-equivariant) and $\alpha^{\vee}$ (the dual of the $\mu_{3}$-action on $P$) on $A$ are inverse to each other. $A$.
Lemma 3.12.
$\operatorname{Prym}(C,\gamma):=(P,\lambda,\alpha)$ and $\operatorname{Prym}(C,\gamma)^{\vee}:=(A,\widehat{\lambda},\hat{\alpha})$ are indecomposable $\mu_{3}$-abelian surfaces over $k$.
Proof.
The first condition of Definition 3.7 for $\operatorname{Prym}(C,\gamma)^{\vee}$ follows from Proposition 2.15 and the fact that $\hat{\alpha}$ preserves $C\hookrightarrow A$. Moreover, since $C$ is irreducible, $(A,\widehat{\lambda})$ is indecomposable. The second condition for $\operatorname{Prym}(C,\gamma)$ follows from Lemma 2.13. The lemma now follows from bigonal duality (Proposition 2.19). ∎
The next theorem is a strong Torelli statement for the association $(C,\gamma)\mapsto\operatorname{Prym}(C,\gamma)^{\vee}$. To state it, we upgrade the set of marked bielliptic Picard curves (respectively $\mu_{3}$-abelian surfaces) to a category, where the morphisms are isomorphisms of curves (respectively polarized abelian surfaces) respecting the $\mu_{6}$-structure (respectively $\mu_{3}$-structure).
Theorem 3.13.
The association $(C,\gamma)\mapsto\operatorname{Prym}(C,\gamma)^{\vee}$ induces an equivalence of categories:
[…]
Proof.
This association is well defined by Lemma 3.12. It is functorial: an isomorphism of marked bielliptic Picard curves $(C,\gamma)\rightarrow(C^{\prime},\gamma^{\prime})$ induces an isomorphism $\operatorname{Prym}(C,\gamma)\rightarrow\operatorname{Prym}(C^{\prime},\gamma^{\prime})$ of marked $\mu_{3}$-abelian surfaces. We explicitly construct a quasi-inverse. To a triple $(A,\lambda,\alpha)$, in the notation of Proposition 3.11, we associate the marked bielliptic Picard curve $(X_{1},\gamma_{1})$. Since the constructions of §3.1 and §3.2 are canonical, this association is also functorial. The fact that this is a quasi-inverse follows from Propositions 2.15 and 3.4. ∎
By Lemma 2.8, the left hand side of (3.7) is in bijection with the set of equivalence classes $\{(a,b)\in k\times k\mid b(a^{2}-4b)\neq 0\}/\sim$, where $(a,b)\sim(a^{\prime},b^{\prime})$ if $(a,b)=(\lambda^{6}a^{\prime},\lambda^{12}b^{\prime})$ for some $\lambda\in k^{\times}$. If $k=k^{\operatorname{sep}}$, the $j$-invariant (2.3) maps this set bijectively to the ‘$j$-line’ $k\setminus\{0\}$. Theorem 3.13 thus shows that this description also applies to the right hand side of (3.7). In particular, we may define the $j$-invariant of a $\mu_{3}$-abelian surface $(A,\lambda,\alpha)$ as $j((C,\gamma))$, where $(C,\gamma)$ is any bielliptic Picard curve with $\operatorname{Prym}(C,\gamma)^{\vee}=(A,\lambda,\alpha)$.
Remark 3.14.
Since $\operatorname{Prym}(C,\gamma)\simeq\operatorname{Prym}(\widehat{C},\hat{\gamma})^{\vee}$ (Proposition 2.19) and bigonal duality induces an autoequivalence on the category of marked bielliptic Picard curves, Theorem 3.13 implies that the association $(C,\gamma)\mapsto\operatorname{Prym}(C,\gamma)$ also induces an equivalence of categories of the form (3.7). On the other hand, the association $(A,\lambda,\alpha)\mapsto(A,\lambda,\alpha^{-1})$ is again an auto-equivalence on the category of marked $\mu_{3}$-abelian surfaces. In Corollary 5.5 we will prove a precise relationship between these equivalences.
Remark 3.15.
Barth has shown that Torelli fails for the family of all bielliptic genus $3$ curves $(X,\tau)$. Suppose that $k$ is algebraically closed of characteristic zero and $A$ is an indecomposable $(1,2)$-polarized abelian surface $(A,\lambda)/k$. Then there exist many bielliptic genus $3$ curves $(X,\tau)$ over $k$ with dual Prym variety isomorphic to $(A,\lambda)$; by Proposition 3.4 any smooth member of the pencil $|\mathscr{L}|$ on $A$ will do. Theorem 3.13 shows that we can correct this when $C$ is a bielliptic Picard curve and we fix the $\mu_{6}$-action on $C$ and the $\mu_{3}$-action on $A$.
Remark 3.16.
Let $S$ be a ${\rm\mathbb{Z}}[1/6]$-scheme. A marked bielliptic Picard curve over $S$ is a smooth proper morphism $C\rightarrow S$ together with a $\mu_{6}$-action $\gamma\colon\mu_{6}\times_{S}C\rightarrow C$ such that for every geometric point $\bar{s}\in S$, $(C_{\bar{s}},\gamma_{\bar{s}})$ is a marked bielliptic Picard curve. With the evident notion of isomorphisms, this defines the stack $\mathfrak{M}$ of (marked) bielliptic Picard curves over ${\rm\mathbb{Z}}[1/6]$ (in the fppf topology), isomorphic to an open substack of the weighted projective stack ${\rm\mathbb{P}}(6,12)$. A $\mu_{3}$-abelian surface over $S$ is a $(1,2)$-polarized abelian scheme $(A,\lambda)$ over $S$ together with a $\mu_{3}$-action $\alpha\colon\mu_{3}\times_{S}A\rightarrow A$ such that for every geometric point $\bar{s}\in S$, $(A_{\bar{s}},\lambda_{\bar{s}},\alpha_{\bar{s}})$ is a $\mu_{3}$-abelian surface. This defines the stack $\mathfrak{Y}$ of $\mu_{3}$-abelian surfaces over ${\rm\mathbb{Z}}[1/6]$. There exists an open substack $\mathfrak{Y}^{\text{indec}}\subset\mathfrak{Y}$ of those surfaces such that each geometric fiber is indecomposable. Theorem 3.13 then generalizes to the statement that the association $(C,\gamma)\mapsto\operatorname{Prym}(C,\gamma)^{\vee}$ induces an isomorphism of stacks $\mathfrak{M}\xrightarrow{\sim}\mathfrak{Y}^{\text{indec}}$.
4. Shimura curves and quaternionic multiplication
We relate the moduli space of bielliptic Picard curves to certain unitary and quaternionic Shimura curves, which will allow us to deduce that these bielliptic Picard Pryms surfaces $P$ have quaternionic multiplication over $\overline{k}$ (Corollary 4.8). This will also help us study which $P$ are geometrically non-simple in §6.4, since these correspond to CM points on the Shimura curve side.
4.1. The quaternion order $\mathcal{O}$
Let $B$ be the quaternion algebra $(-3,2)_{{\rm\mathbb{Q}}}$ with ${\rm\mathbb{Q}}$-basis $\{1,i,j,ij\}$ and relations $i^{2}=-3$, $j^{2}=2$, $ij=-ji$. This is an indefinite quaternion algebra of discriminant $6$. Let $\omega:=\frac{1}{2}(-1+i)$, so that $\omega^{2}+\omega+1=0$. Let $\mathcal{O}:={\rm\mathbb{Z}}[\omega,j]={\rm\mathbb{Z}}+{\rm\mathbb{Z}}\omega+{\rm\mathbb{Z}}j+{\rm\mathbb{Z}}\omega j$. A discriminant calculation shows that $\mathcal{O}$ is a maximal order in $B$. As an associative ring, $\mathcal{O}$ has generators $\omega,j$ and relations $\omega^{2}+\omega+1=0$, $j^{2}=2$ and $j\omega=\omega^{-1}j$.
Let ${\rm Aut}(\mathcal{O})$ denote the set of ring automorphisms of $\mathcal{O}$, which we will think of as acting on the right on $\mathcal{O}$. By the Skolem–Noether theorem, every element of ${\rm Aut}(\mathcal{O})$ is given by conjugating by an element of $b\in B^{\times}$ (uniquely defined up to ${\rm\mathbb{Q}}^{\times}$) normalizing $\mathcal{O}$; denote the conjugation action of such an element $b$ by $[b]\in{\rm Aut}(\mathcal{O})$.
For reasons that will become clear in §6.2, we define the groups
${\rm Aut}_{i}^{+}(\mathcal{O})$ :=$:=\{\varphi\in{\rm Aut}(\mathcal{O})\mid i^{\varphi}=i\},$
${\rm Aut}_{i}(\mathcal{O})$ :=$:=\{\varphi\in{\rm Aut}(\mathcal{O})\mid i^{\varphi}=\pm i\}.$
Lemma 4.1.
${\rm Aut}_{i}^{+}(\mathcal{O})=\langle[1-\omega]\rangle$ is cyclic of order $6$ and ${\rm Aut}_{i}(\mathcal{O})=\langle[1-\omega],[j]\rangle$ is dihedral of order $12$.
Proof.
A direct calculation shows that $\langle[1-\omega],[j]\rangle\subset{\rm Aut}_{i}(\mathcal{O})$, that this subgroup is dihedral of order $12$ and that $\langle[1-\omega],[j]\rangle\cap{\rm Aut}_{i}^{+}(\mathcal{O})=\langle[1-\omega]\rangle$. Conversely, suppose $[b]\in{\rm Aut}_{i}(\mathcal{O})$. After possibly multplying by $j$, we may assume $i^{[b]}=b^{-1}ib=i$. In that case $b$ centralizes $i$, so $b\in{\rm\mathbb{Q}}(i)$. By [28, Lemma 2.2.1], $[b]$ can be represented by an element of $\mathcal{O}\cap{\rm\mathbb{Q}}(i)={\rm\mathbb{Z}}[\omega]$ of norm dividing $6$. Therefore $[b]\in\langle[1-\omega]\rangle$, as desired. ∎
Lemma 4.2.
Let $G\leq{\rm Aut}_{i}(\mathcal{O})$ be a subgroup. Then the subring $\mathcal{O}^{G}$ fixed by $G$ is: $\mathcal{O}$ if $G=\{1\}$; ${\rm\mathbb{Z}}[\omega]$ if $\{1\}\neq G\subset\langle[1-\omega]\rangle$; isomorphic to ${\rm\mathbb{Z}}[\sqrt{2}]$ if $G$ is conjugate to $\langle[j]\rangle$; isomorphic to ${\rm\mathbb{Z}}[\sqrt{6}]$ if $G$ is conjugate to $\langle[ij]\rangle$. In all other cases, $G$ is dihedral of order $6$ or $12$ and $\mathcal{O}^{G}={\rm\mathbb{Z}}$.
Proof.
This is a direct calculation using the explicit description of ${\rm Aut}_{i}(\mathcal{O})$ from Lemma 4.1. ∎
4.2. Moduli of marked $\mu_{3}$-abelian surfaces
In Remark 3.16, we have introduced the stack $\mathfrak{Y}$ of $\mu_{3}$-abelian surfaces $(A,\lambda,\alpha)$ over ${\rm\mathbb{Z}}[1/6]$, which we now view as a stack over ${\rm\mathbb{Q}}$. This is a Deligne–Mumford stack; let $Y/{\rm\mathbb{Q}}$ be its coarse space. By a variant of [26, Proposition 2.1] for $(1,2)$-polarizations, $Y$ is smooth and purely one-dimensional. The next lemma shows that $Y$ may be viewed as a compactification of the $j$-line ${\rm\mathbb{P}}^{1}\setminus\{0,\infty\}$ from §2.3.
Lemma 4.3.
There exists a unique isomorphism $j\colon Y\xrightarrow{\sim}{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}$ with the property that $j([\operatorname{Prym}(C,\gamma)^{\vee}])=j(C)$ for all marked bielliptic Picard curves $(C,\gamma)$ over $\bar{{\rm\mathbb{Q}}}$.
Proof.
The open substack $\mathfrak{Y}^{\text{indec}}\subset\mathfrak{Y}$ parametrizing those triples $(A,\lambda,\alpha)$ for which $(A,\lambda)$ is indecomposable has coarse space an open subset $Y^{\text{indec}}\subset Y$. Every decomposable $\mu_{3}$-abelian surface $(A,\lambda,\alpha)$ over ${\rm\mathbb{C}}$ is isomorphic to $(E\times E,\lambda_{O}\times 2\lambda_{O},\alpha_{1}\times\alpha_{2})$, where $E:y^{2}=x^{3}+1$, $\lambda_{O}$ is the unique principal polarization on $E$ and $\alpha_{i}$ are mutually inverse nontrivial $\mu_{3}$-actions on $E$. Since there are two choices for the pair $(\alpha_{1},\alpha_{2})$, even up to isomorphism, $(Y\setminus Y^{\text{indec}})({\rm\mathbb{C}})$ has size $2$. Since $Y$ is purely one-dimensional, $Y^{\text{indec}}$ is dense in $Y$. Theorem 3.13 and Lemma 2.8 show that the morphism ${\rm\mathbb{A}}^{2}_{{\rm\mathbb{Q}}}\setminus\{\Delta_{a,b}=0\}\rightarrow Y^{\text{indec}},(a,b)\mapsto[\operatorname{Prym}(C_{a,b},\gamma_{a,b})^{\vee}]$ is surjective, factors through the $j$-invariant morphism ${\rm\mathbb{A}}^{2}_{{\rm\mathbb{Q}}}\setminus\{\Delta_{a,b}=0\}\rightarrow{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}\setminus\{0,\infty\},(a,b)\mapsto j_{a,b}$ of (2.3), and induces an isomorphism ${\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}\setminus\{0,\infty\}\simeq Y^{\text{indec}},j\mapsto[\operatorname{Prym}(C_{a,b},\gamma_{a,b})^{\vee}]$, where $a,b$ are any elements with $j=j_{a,b}$. Putting everything together, $Y$ is a purely one-dimensional smooth variety having an open subset isomorphic to ${\rm\mathbb{P}}^{1}\setminus\{0,\infty\}$ whose complement has size $2$. We conclude that $Y\simeq{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}$ and that the isomorphism $j\colon Y^{\text{indec}}\rightarrow{\rm\mathbb{P}}_{{\rm\mathbb{Q}}}^{1}\setminus\{0,\infty\}$ uniquely extends to an isomorphism $Y\rightarrow{\rm\mathbb{P}}_{{\rm\mathbb{Q}}}^{1}$. ∎
Remark 4.4.
The curve $Y_{{\rm\mathbb{C}}}$ is a disjoint union of Shimura varieties for groups of type $\mathrm{GU}(1,1)$. Similarly to [26, Proposition 3.1], the connected components of $Y_{{\rm\mathbb{C}}}$ are parametrized by isometry classes of rank $2$ Hermitian ${\rm\mathbb{Z}}[\omega]$-lattices of discriminant $2$. Lemma 4.3 shows that there is precisely one such Hermitian lattice, although this is certainly not the easiest way to see that.
Remark 4.5.
The isomorphism $j\colon Y\rightarrow{\rm\mathbb{P}}^{1}$ identifies $Y^{\text{indec}}$ with ${\rm\mathbb{P}}^{1}\setminus\{0,\infty\}$. The two decomposable points $0,\infty$ might be considered ‘cusps’.
4.3. Comparison with a quaternionic Shimura curve
We now compare $Y$ to a quaternionic Shimura curve, in the same spirit as [22, §4.2]. However, our analysis is a little different since we must consider twists and non-principal polarizations. We refer the reader to [45, Chapter 43] for background on Shimura curves.
Definition 4.6.
Let $X$ be the (quaternionic) Shimura curve of discriminant $6$ over ${\rm\mathbb{Q}}$.
The nice curve $X$ is the coarse space of the moduli stack of triples $(A,\lambda,\iota)$, where $(A,\lambda)$ is a $(1,2)$-polarized abelian surface and $\iota\colon\mathcal{O}\rightarrow{\rm End}(A)$ is an embedding such that the Rosati involution restricts to $b\mapsto i\bar{b}i^{-1}$, where we recall from §4.1 that $i=1+2\omega$. (The moduli problem is typically formulated for principal polarizations, but the results of [45, §43.6] have analogues for the polarized order $(\mathcal{O},i)$, see [9, §12].) To such a triple $(A,\lambda,\iota)$ we would like to associate a $\mu_{3}$-abelian surface by restricting $\iota$ to ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\simeq\langle\omega\rangle\subset{\rm\mathbb{Z}}[\omega]^{\times}\subset\mathcal{O}^{\times}$. However, this defines a ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$-action on $A$; to get a $\mu_{3}$-action we need to suitably twist the situation.
To this end, recall the action of the Atkin–Lehner group $W=\{1,w_{2},w_{3},w_{6}\}\simeq{\rm\mathbb{Z}}/2{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/2{\rm\mathbb{Z}}$ on $X$: if $b\in\mathcal{O}$ is any element whose (reduced) norm has absolute value $d\in\{2,3,6\}$, then $w_{d}([A,\lambda,\iota])=[A,\lambda,\iota\circ[b]]$, where $[b]\in{\rm Aut}(\mathcal{O})$ denotes conjugation by $b$. The following curves will be important:
  • Let $X/w_{3}$ be the quotient of $X$ by $w_{3}$. The $W$-action on $X$ induces an action of $W/\langle w_{3}\rangle=\{1,\bar{w}\}$ on $X/w_{3}$.
  • Let $(X/w_{3})^{\bar{w}}$ be the twist of $X/w_{3}$ along the involution $\bar{w}$ and extension ${\rm\mathbb{Q}}(\omega)/{\rm\mathbb{Q}}$.
  • For each $d\in\{2,6\}$, let $X^{w_{d}}$ be the twist of $X$ along the involution $w_{d}$ and extension ${\rm\mathbb{Q}}(\omega)/{\rm\mathbb{Q}}$.
We recall that the twist of a nice curve $C/k$ along an involution $\tau$ and quadratic extension $K/k$ corresponds to the cocycle given by the image of the nontrivial element under $\operatorname{H}^{1}({\rm Gal}(K/k),\langle\tau\rangle)\rightarrow\operatorname{H}^{1}(k,\mathbf{Aut}(C))$. Note that $W$ acts on $X^{w_{2}}$ and $X^{w_{6}}$, $\bar{w}$ acts on $(X/w_{3})^{\bar{w}}$, and we have canonical isomorphisms $X^{w_{2}}/w_{3}\simeq X^{w_{6}}/w_{3}\simeq(X/w_{3})^{\bar{w}}$.
To compare $X$ and $Y$, we will construct morphisms $\pi_{d}\colon X^{w_{d}}\rightarrow Y$ for each $d\in\{2,6\}$. In the following paragraph, for a field $k/{\rm\mathbb{Q}}$ we write $K:=k\otimes_{{\rm\mathbb{Q}}}{\rm\mathbb{Q}}(\omega)$ and ${\rm Aut}_{k}(K)=\{1,\sigma\}$. The curve $X^{w_{2}}$ is the coarse space of the moduli stack whose $k$-points parametrizes triples $(A,\lambda,\iota)$, where $(A,\lambda)$ is a $(1,2)$-polarized abelian surface over $k$ and $\iota\colon\mathcal{O}\rightarrow{\rm End}(A_{K})$ is an embedding such that the Rosati involution restricts to $b\mapsto i\bar{b}i^{-1}$ and such that $\iota^{\sigma}=\iota\circ[j]$. We claim that such a triple naturally gives rise to a $\mu_{3}$-abelian surface $(A,\lambda,\iota|_{\langle\omega\rangle})$ over $k$. Indeed, restricting $\iota$ to $\langle\omega\rangle\subset{\rm\mathbb{Z}}[\omega]^{\times}\subset\mathcal{O}^{\times}$ defines homomorphism ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\rightarrow{\rm End}(A_{K})$ which due to the twisting descends to a $\mu_{3}$-action on $A$. The triple $(A,\lambda,\iota|_{\langle\omega\rangle})$ satisfies first condition of Definition 3.7 since the Rosati involution sends $\iota(\omega)$ to $\iota(\omega^{-1})$. The second condition follows from the identity $\omega^{2}+\omega+1=0$. The association $(A,\lambda,\iota)\mapsto(A,\lambda,\iota|_{\langle\omega\rangle})$ also works on the level of ${\rm\mathbb{Q}}$-schemes, hence it induces a morphism $\pi_{2}\colon X^{w_{2}}\rightarrow Y$. Repeating the above discussion but insisting that $\iota^{\sigma}=\iota\circ[ij]$ instead of $\iota^{\sigma}=\iota\circ[j]$, we obtain a morphism $\pi_{6}\colon X^{w_{6}}\rightarrow Y$.
Proposition 4.7.
For each $d\in\{2,6\}$, $\pi_{d}\colon X^{w_{d}}\rightarrow Y$ induces an isomorphism $\phi_{d}\colon X^{w_{d}}/w_{3}\xrightarrow{\sim}Y$. Under the canonical isomorphisms $X^{w_{2}}/w_{3}\simeq X^{w_{6}}/w_{3}\simeq(X/w_{3})^{\bar{w}}$, $\phi_{2}$ and $\phi_{6}$ correspond to the same isomorphism $(X/w_{3})^{\bar{w}}\simeq Y$.
Proof.
It suffices to check these claims over ${\rm\mathbb{C}}$, so let $\pi=\pi_{2,{\rm\mathbb{C}}}=\pi_{6,{\rm\mathbb{C}}}$. The morphism $\pi$ is given by mapping an isomorphism class of triples $[A,\lambda,\iota]$ to $[A,\lambda,\iota|_{\langle\omega\rangle}]$. We have $\pi\circ w_{3}=\pi$, since $w_{3}([A,\lambda,\iota])=[A,\lambda,\iota\circ[i]]$ and $\iota|_{\langle\omega\rangle}=(\iota\circ[i])|_{\langle\omega\rangle}$. (Recall that $i=1+2\omega$.)
Since $\pi$ is a morphism between nice curves, it is either constant or a finite covering map. To prove the proposition, it suffices to show that $\pi$ has degree $2$, equivalently infinitely many fibres $\pi^{-1}(y)$ have size $2$. Let $x=[A,\lambda,\iota]\in X({\rm\mathbb{C}})$ be a non-CM point and $y=\pi(x)=[A,\lambda,\iota|_{\langle\omega\rangle}]$; there are infinitely many (in fact uncountably many) such points $x$ by the theory of Shimura curves [13, §2.4]. The set $\pi^{-1}(y)$ is in bijection with the set of isomorphism classes of triples $(A^{\prime},\lambda^{\prime},\iota^{\prime})$ such that $(A,\lambda,\iota|_{\langle\omega\rangle})\simeq(A^{\prime},\lambda^{\prime},\iota^{\prime}|_{\langle\omega\rangle})$. Replacing $(A^{\prime},\lambda^{\prime},\iota^{\prime})$ by an isomorphic triple, we may assume that $(A^{\prime},\lambda^{\prime})=(A,\lambda)$ and $\iota^{\prime}|_{{\rm\mathbb{Z}}[\omega]}=\iota|_{{\rm\mathbb{Z}}[\omega]}$. Therefore $\pi^{-1}(y)$ is in bijection with the orbit space
(4.1) $\{\iota^{\prime}\colon\mathcal{O}\rightarrow{\rm End}(A)\mid\iota^{\prime}\text{ compatible with Rosati involution and }\iota^{\prime}|_{{\rm\mathbb{Z}}[\omega]}=\iota|_{{\rm\mathbb{Z}}[\omega]}\}/{\rm Aut}(A,\lambda).$
Since $x$ is assumed to be a non-CM point and $\mathcal{O}$ is maximal, ${\rm End}(A)\simeq\mathcal{O}$. It follows that every $\iota^{\prime}$ of (4.1) is of the form $\iota\circ\alpha$ for some $\alpha\in{\rm Aut}(\mathcal{O})$. Since $\iota^{\prime}|_{{\rm\mathbb{Z}}[\omega]}=\iota|_{{\rm\mathbb{Z}}[\omega]}$, $\alpha$ lies in the subgroup ${\rm Aut}_{i}^{+}(\mathcal{O})=\langle[1-\omega]\rangle$ of Lemma 4.1 (using the notation of §4.1). Conversely, every element $[b]\in\langle[1-\omega]\rangle$ defines an element of (4.1), and two elements $[b],[b^{\prime}]$ lie in the same ${\rm Aut}(A,\lambda)$-orbit if and only if $b^{\prime}b^{-1}\in{\rm\mathbb{Q}}^{\times}\mathcal{O}^{\times}$. It follows that this orbit space has size $|\pi^{-1}(y)|=|\langle[1-\omega]\rangle/\langle[\omega]\rangle|=2$, as desired. ∎
Corollary 4.8.
Every $\mu_{3}$-abelian surface $(A,\lambda,\alpha)$ over a field $k$ of characteristic 00 admits an embedding $\mathcal{O}\hookrightarrow{\rm End}(A_{k^{\operatorname{sep}}})$.
Proof.
Proposition 4.7 shows that the point $[A,\lambda,\alpha]\in Y(k)$ lifts to a $k^{\operatorname{sep}}$-point of $X^{w_{2}}$ under $\pi_{2}$. The corollary follows from the moduli interpretation of $X$. ∎
Remark 4.9.
Corollary 4.8 was proved by Petkova and Shiga over ${\rm\mathbb{C}}$ using period matrix calculations [36, Proposition 7.1].
Our proof of Corollary 4.8 shows that there exists an action of $\mathcal{O}$ on $P_{k^{\operatorname{sep}}}$, but it sheds no light on what this action actually is. In Proposition 6.1 we will give a more geometric proof of the quaternionic multiplication of $P_{k^{\operatorname{sep}}}$, which works in all characteristics $\neq 2,3$ and gives finer arithmetic information, such as the field of definition of the quaternionic multiplication, see §6.2.
5. Identifying bielliptic Picard curves in the pencil
Let $(C,\gamma)$ be a marked bielliptic Picard curve. Its Prym variety $\operatorname{Prym}(C,\gamma)=(P,\lambda,\alpha)$ is an indecomposable $\mu_{3}$-abelian surface (Lemma 3.12). Therefore the results of §3.1-3.2 apply, and the pencil of curves on $P$ contains exactly two bielliptic Picard curves by Proposition 3.11. We have already seen in §2.7 that this pencil contains the bigonal dual curve $\widehat{C}$. In this section we will determine the other bielliptic Picard curve in the pencil: it turns out to be a sextic twist of the original curve $C$! This calculation is a crucial ingredient for the results of §6.
Let $\widehat{C}\subset P$ be the dual curve of $C$ from §2.7. Then $\mathscr{M}:=\mathcal{O}_{P}(\widehat{C})$ is the unique line bundle on $P$ representing the $(1,2)$-polarization with base locus $P[\lambda]$ (Proposition 3.2). By Lemma 2.18, $\widehat{C}$ is a bielliptic Picard curve preserved by the $\mu_{3}$-action $\alpha$. By Proposition 3.11, there is exactly one other curve in $|\mathscr{M}|$ that is preserved by the $\mu_{3}$-action; call this curve $C_{\text{twist}}$. We will explicitly describe $C_{\text{twist}}$ using the work of Ikeda [23].
We will use the notations from §2.5. If $p\in E(k)$, let $\Theta_{p}$ be the image of the map ${\rm Sym}^{2}C\rightarrow J$ sending $Q+Q^{\prime}$ to $Q+Q^{\prime}-\pi^{*}(p)$. Then $D_{p}:=\Theta_{p}\cap P$ is a divisor on $P$. Note that $D_{\pi(\infty)}=\widehat{C}$. Recall that we view $E$ as an elliptic curve with origin $O_{E}:=\pi(\infty)\in E(k)$.
Lemma 5.1.
  1. (1)
    For every $p\in E(k)$, $D_{p}\in|\mathscr{M}|(k)$.
  2. (2)
    If $k=k^{\operatorname{sep}}$, every element of $|\mathscr{M}|$ is of the form $D_{p}$ for some $p\in E(k)$.
  3. (3)
    For every $p\in E(k)$, $D_{\iota(p)}=D_{p}$, where $\iota(p)$ denotes the inverse of $p$ under the group law.
Proof.
This is [23, Lemma 3.4] and [23, Remark 3.5]. ∎
Therefore the map $E\rightarrow|\mathscr{M}|,p\mapsto D_{p}$ factors through the map $\phi\colon E\rightarrow|2O_{E}|$ and gives rise to a $\mu_{3}$-equivariant map $|2O_{E}|\rightarrow|\mathscr{M}|,y\mapsto D_{y}$. We note that if $y$ is $k$-rational, then so is $D_{y}$.
By Lemma 2.8, we may and do assume that $(C,\gamma)=(C_{a,b},\gamma_{a,b})$ for some $a,b\in k$. Then $E_{a,b}$ has equation $y^{3}=x^{2}+ax+b$ and $y$ induces an isomorphism $|2O_{E}|\simeq{\rm\mathbb{P}}^{1}$. There are two fixed points for the $\mu_{3}$-action on $|2O_{E}|$, corresponding to $0,\infty\in{\rm\mathbb{P}}^{1}(k)$; call them $y_{0},y_{\infty}\in|2O_{E}|$. Then $D_{y_{\infty}}=D_{O_{E}}=\widehat{C}$ and $D_{y_{0}}=C_{\text{twist}}$.
In order to compute equations for $D_{y_{0}}$, we will give a different description of the family of divisors $\{D_{p}\}_{p\in E}$. For ease of notation, write $C^{(2)}={\rm Sym}^{2}C$. For each $\xi\in{\rm Pic}^{2}(E)$, consider the closed subscheme
$F_{\xi}:=\{Q+Q^{\prime}\in C^{(2)}\mid\pi(Q)+\pi(Q^{\prime})\sim\xi\}\subset C^{(2)}.$
Since $C$ is not hyperelliptic, the map $Q+Q^{\prime}-\pi^{*}(p)\mapsto Q+Q^{\prime}$ induces an isomorphism $\phi_{p}\colon D_{p}\xrightarrow{\sim}F_{2p}$. Consider the morphism
$\psi\colon C^{(2)}$ $\longrightarrow{\rm\mathbb{P}}^{2}$
$([x:y:z],[x^{\prime}:y^{\prime}:z^{\prime}])$ $\mapsto[yz^{\prime}-y^{\prime}z:xz^{\prime}-x^{\prime}z:xy^{\prime}-x^{\prime}y],$
which sends a pair of distinct points on $C$ to the line it spans in the canonical embedding $C\subset{\rm\mathbb{P}}^{2}$. Ikeda has shown that for every $\xi\in{\rm Pic}^{2}(E)$ such that $F_{\xi}$ is smooth, $\psi$ maps $F_{\xi}$ isomorphically onto a plane quartic in ${\rm\mathbb{P}}^{2}$ and determined this quartic explicitly [23, Lemma 3.12]. (In fact, he has shown this for any bielliptic genus $3$ curve.) We will use his calculation to determine $C_{\text{twist}}$, together with a descent argument along a quadratic extension.
Let $\alpha,\alpha^{\prime}$ be the roots of $x^{2}+ax+b$. Let $p=(\alpha,0)$ and $p^{\prime}=(\alpha^{\prime},0)$. Let $K=k(\alpha,\alpha^{\prime})=k(\sqrt{a^{2}-4b})$. Then $p,p^{\prime}\in E(K)$ and $C_{\text{twist}}=D_{y_{0}}=D_{p}=D_{p^{\prime}}$. Both $p$ and $p^{\prime}$ give rise to $K$-isomorphisms $\phi_{p}\colon D_{p}\rightarrow F_{2p}$ and $\phi_{p^{\prime}}\colon D_{p^{\prime}}\rightarrow F_{2p^{\prime}}$. Let $[u:v:1]$ be affine coordinates for the target of the morphism $\psi\colon{\rm Sym}^{2}C\rightarrow{\rm\mathbb{P}}^{2}$.
Lemma 5.2.
$\psi(F_{2p})=\psi(F_{2p^{\prime}})\subset{\rm\mathbb{P}}^{2}_{K}$ is cut out by the projective closure of the equation
(5.1) $(2a^{2}-8b)v^{3}=bu^{4}+au^{2}+1.$
Proof.
This is a special case [23, Lemma 3.12] [^4] The results of loc. cit. assume that $k=\mathbb{C}$, but the proof of the lemma works over any field of characteristic $\neq 2$.. For the reader’s convenience, we provide the translation to Ikeda’s notation. In his notation: $C$ has equation $(z^{2}+\frac{a}{2}y^{2})^{2}=x^{3}y+\left(\frac{1}{4}a^{2}-b\right)y^{4}$, $S(x,y)=\frac{1}{4}a^{2}y^{2}$, $T(x,y)=x^{3}y+(\frac{1}{4}a^{2}-b)y^{4}$, and $p=[0:\alpha:1]$. ∎
[…]
Lemma 5.3.
[…]
Proof.
Let $(u,v)\in\mathcal{Q}_{K}$ and let $Q_{1}+Q_{2}\in F_{2p}$ be the unique element mapping to $(u,v)$ under $\psi|_{F_{2p}}$. Since $p^{\sigma}=p^{\prime}$, a calculation shows that $\chi^{\sigma}\circ\chi^{-1}$ maps $(u,v)$ to $\psi(R_{1}+R_{2})$, where $R_{1}+R_{2}$ is the unique effective divisor on $C$ such that $R_{1}+R_{2}-\pi^{*}(p^{\prime})\sim Q_{1}+Q_{2}-\pi^{*}(p)$. We interpret this geometrically: let $\ell=\psi(Q_{1}+Q_{2})$ be the line spanned by the points $Q_{1},Q_{2}$ in the embedding $C=C_{a,b}\subset{\rm\mathbb{P}}^{2}$. The condition $Q_{1}+Q_{2}\in F_{2p}$ implies that $Q_{1}+Q_{2}+\tau(Q_{1})+\tau(Q_{2})\sim 2\pi^{*}(p)$, where we recall that $\tau\colon C\rightarrow C$ denotes the bielliptic involution. Combining the above two linear equivalences shows that $\tau(Q_{1})+\tau(Q_{2})+R_{1}+R_{2}\sim\pi^{*}(p)+\pi^{*}(p^{\prime})\sim 4\infty$. Since $4\infty$ is canonical, it follows that $\tau(Q_{1}),\tau(Q_{2}),R_{1},R_{2}$ are collinear and $\psi(R_{1}+R_{2})=\psi(\tau(Q_{1})+\tau(Q_{2}))=\tau(\ell)$. In the coordinates $(u,v)$, this shows that $\psi(R_{1}+R_{2})=(-u,v)$, proving the lemma. ∎
Theorem 5.4.
[…]
Proof.
By Lemma 5.2, $\mathcal{Q}\simeq C_{da,d^{2}b}$, where $d=16b(a^{2}-4b)^{4}$. Suppose first that $K=k$. Then the isomorphism $\psi|_{F_{2p}}\circ\phi_{p}\colon C_{\text{twist}}\rightarrow\mathcal{Q}$ are defined over $k$, hence $C_{\text{twist}}\simeq C_{da,d^{2}b}$ too. Since $K=k(\sqrt{a^{2}-4b})$, $a^{2}-4b$ is a square in $k^{\times}$. Therefore $d$ and $\delta$ have the same image in $k^{\times}/k^{\times 6}$, so $C_{\text{twist}}\simeq C_{da,d^{2}b}\simeq C_{\delta a,\delta^{2}b}$ by Lemma 2.8.
Suppose now that $K/k$ is a quadratic extension and let $\sigma\in{\rm Gal}(K/k)$ be the nontrivial element. Lemma 5.3 shows that $C_{\text{twist}}$ is isomorphic to the quadratic twist of $\mathcal{Q}$ along the extension $K/k$ and involution $(u,v)\mapsto(-u,v)$. A computation (using §2.4) shows that this twist is exactly $C_{\delta a,\delta^{2}b}$. ∎
Theorem 5.4 has consequences for the Prym variety of $C$ and its dual. If $(X,\mu,\beta)$ is a $\mu_{3}$-abelian surface in the sense of Definition 3.7, the $\mu_{3}$-action $\beta$ and $\mu_{2}$-action $[-1]$ combine to a $\mu_{6}$-action. Therefore, just as in §2.4, every $\delta\in k^{\times}/k^{\times 6}$ gives rise to a sextic twist $(X_{\delta},\mu_{\delta},\beta_{\delta})=(X,\mu,\beta)_{\delta}$ of $(X,\mu,\beta)$. The formation of the $\mu_{3}$-abelian surfaces $\operatorname{Prym}(C,\gamma)$ and $\operatorname{Prym}(C,\gamma)^{\vee}$ from §3.3 is compatible with sextic twisting: $\operatorname{Prym}((C,\gamma)_{\delta})\simeq\operatorname{Prym}(C,\gamma)_{\delta}$ and $\operatorname{Prym}((C,\gamma)_{\delta})^{\vee}\simeq\operatorname{Prym}(C,\gamma)^{\vee}_{\delta}$.
Corollary 5.5.
Let $(C,\gamma)=(C_{a,b},\gamma_{a,b})$ be a marked bielliptic Picard curve over $k$. Let $\delta=16b(a^{2}-4b)$. Let $\operatorname{Prym}(C,\gamma)=(P,\lambda,\alpha)$ and $\operatorname{Prym}(C,\gamma)^{\vee}=(A,\widehat{\lambda},\hat{\alpha})$. Then $(P,\lambda,\alpha)\simeq(A,\widehat{\lambda},\hat{\alpha}^{-1})_{\delta}$. In particular, $P_{k^{\operatorname{sep}}}$ and $A_{k^{\operatorname{sep}}}$ are isomorphic as $(1,2)$-polarized abelian surfaces and $P_{k^{\operatorname{sep}}}$ is self-dual.
Proof.
The $\mu_{3}$-action $\alpha$ and the involution $[-1]$ on $P$ combine to a $\mu_{6}$-action that restricts to $\mu_{6}$-actions $\hat{\gamma}$ and $\gamma_{\text{twist}}$ on $\widehat{C}$ and $C_{\text{twist}}$ respectively. Lemma 2.18 shows that $(\widehat{C},\hat{\gamma})$ is a marked bielliptic Picard curve; so is $(C_{\text{twist}},\gamma_{\text{twist}}^{-1})$ by Proposition 3.11. Therefore $(P,\lambda,\alpha)\simeq\operatorname{Prym}(\widehat{C},\hat{\gamma})^{\vee}$ and $(P,\lambda,\alpha^{-1})\simeq\operatorname{Prym}(C_{\text{twist}},\gamma_{\text{twist}}^{-1})^{\vee}$ by Proposition 3.4.
It remains to show that $(P,\lambda,\alpha)\simeq(A,\widehat{\lambda},\hat{\alpha}^{-1})_{\delta}$, or equivalently that $\operatorname{Prym}(C_{\text{twist}},\gamma_{\text{twist}}^{-1})^{\vee}\simeq(A,\widehat{\lambda},\hat{\alpha})_{\delta}$. Recall that $(C_{\delta},\gamma_{\delta})$ denotes the sextic twist of $C$ along $\delta$. Theorem 5.4 shows that $(C_{\text{twist}},\gamma_{\text{twist}}^{-1})\simeq(C_{\delta},\gamma_{\delta})$. Therefore
$\operatorname{Prym}(C_{\text{twist}},\gamma_{\text{twist}}^{-1})^{\vee}\simeq\operatorname{Prym}(C_{\delta},\gamma_{\delta})^{\vee}\simeq(A,\hat{\lambda},\hat{\alpha})_{\delta},$
as desired. ∎
Remark 5.6.
In contrast to the results of §2.6, 2.7, 3.1, Corollary 5.5 crucially uses properties specific to bielliptic Picard curves. Indeed, the last sentence of the corollary might fail for a general bielliptic genus $3$ curve.
Remark 5.7.
The isomorphism $A\simeq P_{\delta^{-1}}$ of Corollary 5.5 can be chosen ‘universally’. More precisely, let $R:={\rm\mathbb{Z}}[1/6,a,b,\Delta_{a,b}^{-1}]$ and $S:={\rm Spec}(R)$. Equation (2.1) defines the universal marked bielliptic Picard curve (in the sense of Remark 3.16) with Prym variety $\mathcal{P}\rightarrow S$ and dual $\mathcal{A}\rightarrow S$. Then there exists an isomorphism $\Phi\colon\mathcal{A}\rightarrow\mathcal{P}_{\delta^{-1}}$ that restricts to an isomorphism of Corollary 5.5 for every geometric point of $S$. (This can be seen, for example, by spreading out an isomorphism over the generic point to all of $S$ using [15, §2, Lemma 1].)
As a first application of Corollary 5.5, we show that the Atkin–Lehner involution on $X/w_{3}$ corresponds to inversion on the $j$-line $Y\simeq{\rm\mathbb{P}}^{1}_{j}$.
Lemma 5.8.
Under the isomorphisms $(X/w_{3})^{\bar{w}}\simeq Y\simeq{\rm\mathbb{P}}^{1}_{j}$ of Proposition 4.7 and Lemma 4.3, the involution $\bar{w}$ on $(X/w_{3})^{\bar{w}}$ corresponds to the involution $j\mapsto 1/j$ on ${\rm\mathbb{P}}^{1}_{j}$.
Proof.
We use the notations of §4.2, 4.3. It suffices to check the claim over ${\rm\mathbb{C}}$ on the locus of indecomposable $\mu_{3}$-abelian surfaces $Y^{\text{indec}}\subset Y$, so let $y=[A,\lambda,\alpha]\in Y^{\text{indec}}({\rm\mathbb{C}})$ be such a point. Let $x=[A,\lambda,\iota]\in X^{w_{2}}({\rm\mathbb{C}})=X({\rm\mathbb{C}})$ be a lift of $y$ under $\pi_{2}$, so $\alpha=\iota|_{\langle\omega\rangle}$. Then $\bar{w}(y)=\pi_{2}(w_{2}(x))$. Since $j$ has norm $2$, $w_{2}(x)=[A,\lambda,\iota\circ[j]]$. Moreover $[j]$ maps $\omega$ to $\omega^{-1}$, so $\iota\circ[j]$ restricts to $\alpha^{-1}$. It follows that $\pi_{2}(w_{2}(x))=[A,\lambda,\alpha^{-1}]$. On the other hand, if $(A,\lambda,\alpha)\simeq\operatorname{Prym}(C,\gamma)^{\vee}$, then Corollary 5.5 shows that $(A,\lambda,\alpha^{-1})\simeq\operatorname{Prym}(C,\gamma)$, which is itself isomorphic to $\operatorname{Prym}(\widehat{C},\hat{\gamma})^{\vee}$, where $(\widehat{C},\hat{\gamma})$ is the bigonal dual curve from §2.7. Since the $j$-invariant of $\widehat{C}$ equals the inverse of the $j$-invariant of $C$ (Lemma 2.20), the lemma follows from the explicit description of the isomorphism $Y\simeq{\rm\mathbb{P}}^{1}_{j}$. ∎
6. Explicit quaternionic multiplication
In Corollary 4.8 we proved that there exists an action of $\mathcal{O}$ on $P_{k^{\operatorname{sep}}}$. Next we use Theorem 5.4 to describe this action explicitly. For the remainder of this paper, we fix a universal isomorphism $\Phi\colon\mathcal{A}\rightarrow\mathcal{P}_{\delta^{-1}}$ as in Remark 5.7.
6.1. Explicit quaternionic multiplication
Let $(C,\gamma)=(C_{a,b},\gamma_{a,b})$ be a marked bielliptic Picard curve over $k$. Let $\delta=\Delta_{a,b}=16b(a^{2}-4b)$, which is nonzero since $C$ is smooth. Consider the ${\rm Gal}_{k}$-set:
(6.1) $\mathcal{S}:=\{(\zeta,\varepsilon)\in k^{\operatorname{sep}}\times k^{\operatorname{sep}}\mid\zeta\text{ is a primitive sixth root of unity and }\varepsilon^{6}=\delta\}.$
Consider $\mu_{3}$-abelian surfaces $\operatorname{Prym}(C,\gamma)=(P,\lambda,\alpha)$ and $\operatorname{Prym}(C,\gamma)^{\vee}=(A,\widehat{\lambda},\hat{\alpha})$. For every $(\zeta,\varepsilon)\in\mathcal{S}$, we will define elements $r_{\zeta},s_{\varepsilon}\in{\rm End}(P_{k^{\operatorname{sep}}})$. Since $\zeta^{2}$ is a third root of unity, we may set $r_{\zeta}:=\alpha(\zeta^{2})\in{\rm End}(P_{k^{\operatorname{sep}}})$.
To define $s_{\varepsilon}$, we use Corollary 5.5. By that corollary and Remark 5.7, the specialization of the universal isomorphism $\Phi$ at the $k$-point $(a,b)\in S(k)$ is an isomorphism $\phi\colon(A,\widehat{\lambda},\hat{\alpha}^{-1})\xrightarrow{\sim}(P,\lambda,\alpha)_{\delta^{-1}}$. The element $\varepsilon$ determines an isomorphism $C_{\delta^{-1},k^{\operatorname{sep}}}\rightarrow C_{k^{\operatorname{sep}}},(x,y)\mapsto(\varepsilon^{3}x,\varepsilon^{4}y)$, hence an isomorphism $\chi_{\varepsilon}\colon(P,\lambda,\alpha)_{\delta^{-1},{k^{\operatorname{sep}}}}\rightarrow(P,\lambda,\alpha)_{k^{\operatorname{sep}}}$ between their Prym varieties. Let $s_{\varepsilon}$ be the composition:
[…]
Proposition 6.1.
For every $x=(\zeta,\varepsilon)\in\mathcal{S}$ the assignment $\omega\mapsto r_{\zeta},j\mapsto s_{\varepsilon}$ extends to an embedding
(6.3) $\iota_{x}\colon\mathcal{O}\rightarrow{\rm End}(P_{k^{\operatorname{sep}}})={\rm End}(P_{\bar{k}}).$
Proof.
We may assume $k=k^{\operatorname{sep}}$. Fix $x=(\zeta,\varepsilon)\in\mathcal{S}$ and write $r=r_{\zeta}$, $s=s_{\varepsilon}$. We will show that $r^{2}+r+1=0$, $s^{2}=2$ and $sr=r^{-1}s$. Since the associative ring $\mathcal{O}$ from §4.1 is generated by $\omega,j$ and satisfies the same three relations, this will imply that there exists a unique ring homomorphism $\iota_{x}\colon\mathcal{O}\rightarrow{\rm End}(P_{k^{\operatorname{sep}}})$ with $\iota_{x}(\omega)=r$ and $\iota_{x}(j)=s$. Such a homomorphism is automatically an embedding since $\mathcal{O}\otimes{\rm\mathbb{Q}}$ is a simple algebra.
Equation (2.4) of Lemma 2.14 shows that $r^{2}+r+1=0$. To check that $sr=r^{-1}s$, note that the surfaces $P,P_{\delta^{-1}},A$ have $\mu_{3}$-actions $\alpha,\alpha_{\delta^{-1}},\hat{\alpha}$ respectively. The morphisms $\lambda$ and $\chi_{\varepsilon}$ from (6.2) are equivariant with respect to these $\mu_{3}$-actions. On the other hand, $\phi$ intertwines the $\mu_{3}$-action on the domain with the inverse of the $\mu_{3}$-action on the target. The same must therefore be true for their composition $s$, so $sr=r^{-1}s$.
It suffices to prove that $s^{2}=2$; we will use the results of §4.3 to achieve this. Write $\psi=\chi_{\varepsilon}\circ\phi$. The morphism $\psi$ is compatible with the polarizations since $\phi$ and $\chi_{\varepsilon}$ are, so $\psi^{\vee}\circ\lambda\circ\psi=\widehat{\lambda}$. Therefore
(6.4) $s^{2}=\psi\lambda\psi\lambda=(\psi(\psi^{\vee})^{-1})(\psi^{\vee}\lambda\psi\lambda)=(\psi(\psi^{\vee})^{-1})(\widehat{\lambda}\lambda)=2(\psi(\psi^{\vee})^{-1}).$
It follows that $s^{2}=2$ if and only if $\psi^{\vee}=\psi$. To prove the latter statement, we will first show that $(\psi^{\vee})^{-1}\circ\psi\in{\rm Aut}(A,\widehat{\lambda},\hat{\alpha})$. The identity $\psi^{\vee}\circ\lambda=\widehat{\lambda}\circ\psi$ implies that $\lambda\circ(\psi\circ\lambda\circ\psi^{\vee}-\widehat{\lambda})\circ\lambda=0$. Since $\lambda$ is an isogeny, the inner expression must be zero, so $\psi\circ\lambda\circ\psi^{\vee}=\widehat{\lambda}$. In other words $\psi^{\vee}$ is compatible with the polarizations. Moreover, $\psi$ intertwines the $\mu_{3}$-action $\hat{\alpha}$ on $A$ with the inverse of the $\mu_{3}$-action $\alpha$ on $P$: $\psi(\hat{\alpha}(\omega)(x))=\alpha(\omega)^{-1}(\psi(x))$ for all $x\in A$ and $\omega\in\mu_{3}$. Applying duality to this identity and using that $\alpha^{\vee}(\omega)=\hat{\alpha}(\omega^{-1})$, $\psi^{\vee}$ again intertwines the $\mu_{3}$-action with the inverse of the other. We conclude that $(\psi^{\vee})^{-1}\circ\psi\in{\rm Aut}(A,\lambda,\alpha)$.
To show that this automorphism is the identity, we will use that its construction can be carried out in families. More precisely, let $R={\rm\mathbb{Z}}[1/6,a,b,\Delta_{a,b}^{-1}]$ and $\tilde{R}=R[\varepsilon]/(\varepsilon^{6}-\Delta_{a,b})$. Equation (2.1) defines a universal marked bielliptic Picard curve $(\mathcal{C},\gamma)$ over $S={\rm Spec}(R)\subset{\rm\mathbb{A}}^{2}_{{\rm\mathbb{Z}}[1/6]}$ with Prym variety $(\mathcal{P},\lambda,\alpha)$ and dual $(\mathcal{A},\widehat{\lambda},\hat{\alpha})$. On the cover $\tilde{S}={\rm Spec}(\tilde{R})\rightarrow S$ given by adjoining a sixth root of $\Delta_{a,b}$, there exists an isomorphism $\Psi\colon\mathcal{A}_{\tilde{S}}\rightarrow\mathcal{P}_{\tilde{S}}$ with the property that $\Psi$ specializes to $\psi$ for all morphisms ${\rm Spec}(k)\rightarrow\tilde{S}$ and that $(\Psi^{\vee})^{-1}\circ\Psi\in{\rm Aut}_{\tilde{S}}(\mathcal{A},\lambda,\alpha)$. Since the moduli stack of $(1,2)$-polarized abelian surfaces is separated and Deligne–Mumford, $\mathbf{Aut}_{\tilde{S}}(\mathcal{A},\lambda,\alpha)\rightarrow\tilde{S}$ is finite and unramified. Since finite unramified morphisms are étale locally on the target disjoint unions of closed immersions [44, Tag 04HJ] and the base $\tilde{S}$ is connected, it suffices to show that $(\Psi^{\vee})^{-1}\Psi$ is the identity for a single specialization $s\in\tilde{S}$.
The specialization $(a_{0},b_{0},\varepsilon_{0})=(1,1/8,1)$ defines a ${\rm\mathbb{Q}}$-point $s_{0}$ of $\tilde{S}$. Theorem 3.13 and Lemma 2.10 shows that $\mathbf{Aut}((\mathcal{A},\lambda,\alpha)_{s_{0}})=\mu_{6}$. Since $\mu_{6}({\rm\mathbb{Q}})=\{\pm 1\}$, either $(\Psi^{\vee})^{-1}\Psi=1$ or $(\Psi^{\vee})^{-1}\Psi=-1$. Assume for the sake of contradiction that $(\Psi^{\vee})^{-1}\Psi=-1$. Then (6.4) shows that $s^{2}=-2$. This implies that $r$ and $s$ determine an embedding of the definite quaternion algebra $(-3,-2)_{{\rm\mathbb{Q}}}$ into ${\rm End}(\mathcal{P}_{s})\otimes{\rm\mathbb{Q}}$ for every specialization $s\in\tilde{S}({\rm\mathbb{C}})$. By the classification of endomorphism algebras of complex abelian surfaces [5, Proposition 5.5.7, Exercise 9.10(1) and Exercise 9.10(4)], this implies that $\mathcal{P}_{s}$ is isogenous to the square of a CM elliptic curve. This is a contradiction, since there are (uncountably) many specializations $s$ for which $\mathcal{P}_{s}$ is simple, using Proposition 4.7 and CM theory for the quaternionic Shimura curve $X$. We conclude that $\psi^{\vee}=\psi$ and hence by (6.4) that $s^{2}=2$, as desired. ∎
If $P_{\bar{k}}$ is simple, then $\mathcal{O}$ is the full ring of endomorphisms of $P_{\bar{k}}$:
Lemma 6.2.
Let $C/k$ be a bielliptic Picard curve with Prym variety $P$ and suppose that $k$ is algebraically closed. Then $P$ is not simple normal-$\Leftrightarrow$ $\mathcal{O}\not\simeq{\rm End}(P)$ normal-$\Leftrightarrow$ $P$ is isogenous to the square of a CM elliptic curve. Consequently, if $P$ is simple then ${\rm End}(P)=\iota_{x}(\mathcal{O})$ for every $x\in\mathcal{S}$.
Proof.
This is a direct consequence of the Albert classification [5, Proposition 5.5.7, Exercise 9.10(1) and Exercise 9.10(4)]. ∎
Remark 6.3.
If $\mathrm{char}(k)=0$ or if $k=\overline{{\rm\mathbb{F}}}_{p}$ with $p>3$, then these conditions are also equivalent to the condition that $P$ is isomorphic to a product of elliptic curves. See [40] for the former case and [38, Theorem 3.1] for the latter.
We will study geometrically split Prym varieties in more detail in §6.4.
6.2. Galois action on the endomorphisms
Keep the notations of the previous section. We proceed to determine how the various embeddings $\iota_{x}\colon\mathcal{O}\rightarrow{\rm End}(P_{k^{\operatorname{sep}}})$ are related and deduce an explicit description of the Galois action on $\mathcal{O}$.
Let ${\rm Sym}(\mathcal{S})$ be the group of bijections $\mathcal{S}\rightarrow\mathcal{S}$ acting on the right on the ${\rm Gal}_{k}$-set $\mathcal{S}$ of (6.1). Define the subgroup $D_{6}\subset{\rm Sym}(\mathcal{S})$ generated by $\rho,\sigma$, where $(\zeta,\varepsilon)^{\rho}=(\zeta,\zeta^{-1}\varepsilon)$ and $(\zeta,\varepsilon)^{\tau}\mapsto(\zeta^{-1},\varepsilon)$. Then $D_{6}$ is a dihedral group of order $12$ and $D_{6}$ acts simply transitively on $\mathcal{S}$.
Recall the subgroup ${\rm Aut}_{i}(\mathcal{O})$ from Lemma 4.1. The assignment $\rho\mapsto[1-\omega]$, $\tau\mapsto[j]$ induces an isomorphism $\varphi\colon D_{6}\xrightarrow{\sim}{\rm Aut}_{i}(\mathcal{O})$.
Lemma 6.4.
For all $x\in\mathcal{S}$, $g\in D_{6}$ and $b\in\mathcal{O}$, $\iota_{x^{g}}(b^{\varphi(g)})=\iota_{x}(b)$
Proof.
It suffices to prove the claimed identity when $g\in\{\tau,\rho\}$ and $b\in\{\omega,j\}$, for which it can be checked using routine but intricate calculations. Write $x=(\zeta,\varepsilon)$. Suppose first that $g=\tau$, then $\iota_{x^{g}}(b^{\varphi(g)})=\iota_{(\zeta^{-1},\varepsilon)}(j^{-1}bj)$. If $b=\omega$, this equals $\iota_{(\zeta^{-1},\varepsilon)}(j^{-1}\omega j)=\iota_{(\zeta^{-1},\varepsilon)}(\omega^{-1})=r_{\zeta^{-1}}^{-1}=r_{\zeta}=\iota_{x}(\omega)$. If $b=j$, this equals $\iota_{(\zeta^{-1},\varepsilon)}(j)=s_{\varepsilon}=\iota_{x}(j)$.
Suppose next that $g=\rho$. If $b=\omega$, both sides are simply $r_{\zeta}$. If $b=j$, then
$\iota_{x^{g}}(b^{\varphi(g)})=\iota_{x^{g}}(j^{[1-\omega]})=\iota_{x^{g}}(-\omega^{-1}j)=-\iota_{(\zeta,\zeta^{-1}\varepsilon)}(\omega^{-1})\iota_{(\zeta,\zeta^{-1}\varepsilon)}(j)=-r_{\zeta}^{-1}s_{\zeta^{-1}\varepsilon}.$
The definition of $s_{\varepsilon}$ in (6.2) shows that $s_{\zeta^{-1}\varepsilon}$ is given by postcomposing $s_{\varepsilon}$ with the automorphism of $P$ induced by $\gamma(\zeta^{-1})$. A simple calculation shows that this is exactly $-r_{\zeta}$, so $s_{\zeta^{-1}\varepsilon}=-r_{\zeta}s_{\varepsilon}$. Therefore $\iota_{x^{g}}(j^{g})=-r_{\zeta}^{-1}(-r_{\zeta}s_{\varepsilon})=s_{\varepsilon}=\iota_{x}(b)$. ∎
[…]
Theorem 6.5.
For all $x\in\mathcal{S}$, $\sigma\in{\rm Gal}_{k}$ and $b\in\mathcal{O}$, $\iota_{x}(b)^{\sigma}=\iota_{x}(b^{\varphi(\sigma^{-1})})$. Consequently, there exists an embedding $\mathcal{O}\subset{\rm End}(P_{k^{\operatorname{sep}}})$ such that $f^{\sigma}=f^{\varphi(\sigma^{-1})}$ for all $f\in\mathcal{O}$ and $\sigma\in{\rm Gal}_{k}$. In particular, $\mathcal{O}$ is ${\rm Gal}_{k}$-stable and defined over $L=k(\omega,\sqrt[6]{\delta})$.
Proof.
Write $x=(\zeta,\varepsilon)$. The explicit construction of $r_{\zeta}$ and $s_{\varepsilon}$ shows that $r_{\zeta}^{\sigma}=r_{\zeta^{\sigma}}$ and $s_{\varepsilon}^{\sigma}=s_{\varepsilon^{\sigma}}$. Therefore $\iota_{x}(b)^{\sigma}=\iota_{x^{\sigma}}(b)$. By the previous lemma, $\iota_{x}(b)^{\sigma}=\iota_{x^{\sigma}}(b)=\iota_{x}(b^{\varphi(\sigma^{-1})})$. ∎
The endomorphism field of an abelian variety is the smallest field extension over which all endomorphisms are defined [20]. If $P$ is geometrically simple, then ${\rm End}(P_{\bar{k}})=\mathcal{O}$ by Lemma 6.2. Since the action of $D_{6}$ on $\mathcal{O}$ is faithful, Theorem 6.5 has the following immediate corollary.
Corollary 6.6.
Let $C=C_{a,b}$ be a bielliptic Picard curve over $k$. Suppose that $P_{\bar{k}}$ is simple. Then the endomorphism field of $P/k$ is $k(\omega,\sqrt[6]{16b(a^{2}-4b)})$.
The endomorphism field is invariant under isogeny, so the corollary holds verbatim for the dual Prym $A/k$.
Example 6.7.
Corollary 6.6 gives many examples of abelian surfaces with large endomorphism fields. For example, for the curve $C=C_{3,4}:y^{3}=x^{4}+3x^{2}+4$ over ${\rm\mathbb{Q}}$, the Prym $P=P_{3,4}$ is geometrically simple (see Proposition 6.18) and its endomorphism field is ${\rm\mathbb{Q}}(\sqrt{-3},\sqrt[6]{-7})$, a dihedral degree $12$ extension of ${\rm\mathbb{Q}}$. In the notation of [17], this seems to be the first published example of a geometrically simple abelian surface over ${\rm\mathbb{Q}}$ with Sato–Tate group $J(E_{6})$. Similarly, $P_{-4,2}$ is a geometrically simple abelian surface with endomorphism field ${\rm\mathbb{Q}}(\sqrt{-3},\sqrt[3]{2})$ and Sato–Tate group $J(E_{3})$.
We also record a simple description of the Artin representations ${\rm End}^{0}(P_{k^{\mathrm{sep}}})$ and ${\rm NS}(P_{k^{\mathrm{sep}}})\otimes{\rm\mathbb{Q}}$. Recall that ${\rm NS}(P_{k^{\mathrm{sep}}})$ is the group of line bundles on $P_{k^{\mathrm{sep}}}$ modulo algebraic equivalence. Write $G={\rm Gal}(L/k)$ and view it as a subgroup of $D_{6}$. Let $\chi\colon D_{6}\rightarrow\{\pm 1\}$ be the unique homomorphism with kernel $C_{6}=\langle\rho\rangle$ and let $\mathrm{std}\colon D_{6}\rightarrow{\rm GL}_{2}({\rm\mathbb{Q}})$ be a model for the unique $2$-dimensional irreducible faithful representation of $D_{6}$.
Corollary 6.8.
Suppose that $P_{\bar{k}}$ is simple. Then there are isomorphisms of $G$-representations ${\rm End}^{0}(P_{k^{\mathrm{sep}}})\simeq(1\oplus\chi\oplus\mathrm{std})|_{G}$ and ${\rm NS}(P_{k^{\mathrm{sep}}})\otimes{\rm\mathbb{Q}}\simeq(1\oplus\mathrm{std})|_{G}$.
Proof.
The first part immediately follows from Theorem 6.5: $1\in B$ generates the trivial representation, $i\in B$ generates $\chi$ and $\{j,ij\}$ generate $\mathrm{std}$. To prove the second isomorphism, recall that the ${\rm Gal}_{k}$-equivariant map $L\mapsto\lambda^{-1}\phi_{L}$ gives an isomorphism between ${\rm NS}(P_{k^{\mathrm{sep}}})\otimes{\rm\mathbb{Q}}$ and the subgroup of ${\rm End}^{0}(P_{k^{\mathrm{sep}}})$ fixed by the Rosati involution. The Rosati involution on ${\rm End}^{0}(P_{k^{\mathrm{sep}}})\simeq B$ is given by $b\mapsto i\bar{b}i^{-1}$, the unique positive involution on $B$ that induces complex conjugation on ${\rm\mathbb{Q}}(i)={\rm\mathbb{Q}}(\omega)$. It follows that ${\rm NS}(P_{k^{\mathrm{sep}}})\otimes{\rm\mathbb{Q}}\simeq\{f\in B\mid i\bar{f}i^{-1}=f\}=\text{span}\{1\}\oplus\text{span}\{j,ij\}\simeq(1\oplus\mathrm{std})|_{G}$. ∎
6.3. Prym surfaces of ${\rm GL}_{2}$-type
Let $C=C_{a,b}$ be a bielliptic Picard curve over $k$ with Prym variety $P$, and let $\iota\colon\mathcal{O}\hookrightarrow{\rm End}(P_{k^{\operatorname{sep}}})$ be an embedding satisfying the conclusion of Theorem 6.5. Recall that $k(\omega)$ denotes the smallest field extension of $k$ containing a primitive third root of unity.
Corollary 6.9.
The following conditions are equivalent:
  1. (1)
    $16b(a^{2}-4b)$ is a sixth power in $k(\omega)$;
  2. (2)
    $\iota(\mathcal{O})\subset{\rm End}(P_{k(\omega)})$;
  3. (3)
    $\iota(\mathcal{O})\cap{\rm End}(P)$ is either an order in a real quadratic field or all of $\iota(\mathcal{O})$.
Proof.
$(1)$ and $(2)$ are equivalent by Theorem 6.5. To prove the equivalence between $(2)$ and $(3)$, let $G$ be the image of the homomorphism $\varphi\colon{\rm Gal}_{k}\rightarrow{\rm Aut}_{i}(\mathcal{O})$ of Theorem 6.5, and let $H=\text{Im}({\rm Gal}_{k(\omega)}\rightarrow{\rm Aut}_{i}^{+}(\mathcal{O}))$. Then $H=G\cap{\rm Aut}_{i}^{+}(\mathcal{O})$, where ${\rm Aut}_{i}^{+}(\mathcal{O})=\langle[1-\omega]\rangle$. A group theory calculation using Lemma 4.2 shows that $H=\{1\}$ (in other words, $(2)$ holds) if and only if $\mathcal{O}^{G}$ contains an order in a real quadratic field (in other words $(3)$ holds). ∎
Remark 6.10.
Recall that an abelian surface $X/k$ is of ${\rm GL}_{2}$-type if ${\rm End}(X)$ contains a quadratic ring. If $P$ is geometrically simple and $k\neq k(\omega)$ then $\iota(\mathcal{O})\cap{\rm End}(P)={\rm End}(P)$ (Lemma 6.2) and ${\rm End}(P)$ is either ${\rm\mathbb{Z}}$ or an order in a real quadratic field by Lemma 4.2, so the conditions above are equivalent to $P$ being of ${\rm GL}_{2}$-type. If $k=k(\omega)$ then ${\rm\mathbb{Z}}[\omega]\hookrightarrow{\rm End}(P)$, so $P$ is always of ${\rm GL}_{2}$-type.
We can be more precise about the ring of endomorphisms ${\rm End}(P)$. A calculation shows that an element $\delta\in k^{\times}$ is a sixth power in $k(\omega)$ if and only if either $\delta$ or $-27\delta$ is a sixth power in $k$. This dichotomy breaks up Corollary 6.9 into the following two cases.
Corollary 6.11.
Assume that $j(C)\neq 1$. Then the following conditions are equivalent.
  1. (1)
    $\Delta_{a,b}=16b(a^{2}-4b)$ is a sixth power in $k$.
  2. (2)
    ${\rm End}(P)\cap\iota(\mathcal{O})$ contains a subring isomorphic to ${\rm\mathbb{Z}}[\sqrt{2}]$.
  3. (3)
    There exist $t,d\in k^{\times}$ such that $(a,b)=(2(t^{2}+1)^{2}td^{3},(t^{2}+1)^{3}t^{2}d^{6})$.
If these conditions are satisfied then $P$ and $A=P^{\vee}$ are $k$-isomorphic. A sextic twist of $P$ satisfies the above conditions if and only if $j(C)=-t^{2}$ for some $t\in k^{\times}$.
Proof.
$(1)$ implies $P\simeq A$ by Corollary 5.5. To show $(1)\Leftrightarrow(2)$, let $H\leq G$ be the subgroups of ${\rm Aut}_{i}(\mathcal{O})$ defined in the proof of Corollary 6.9. Then $\mathcal{O}^{G}$ contains a subring isomorphic to ${\rm\mathbb{Z}}[\sqrt{2}]$ if and only if $G$ is conjugate to a subgroup of $\langle[j]\rangle$ by Lemma 4.2. Theorem 6.5 and Galois theory show that this is equivalent to $T^{6}-\delta$ having a $k$-rational root. This proves that $(1)\Leftrightarrow(2)$. The equivalence of $(1)$ and $(3)$ as well as the last claim of the corollary are routine algebra. ∎
Corollary 6.12.
Assume that $j(C)\neq 1$. Then the following conditions are equivalent.
  1. (1)
    $-27\Delta_{a,b}=-432b(a^{2}-4b)$ is a sixth power in $k$.
  2. (2)
    ${\rm End}(P)\cap\iota(\mathcal{O})$ contains a subring isomorphic to ${\rm\mathbb{Z}}[\sqrt{6}]$.
  3. (3)
    There exist $t,d\in k^{\times}$ such that $(a,b)=(18d^{3}t(1-3t^{2})^{2},3^{4}d^{6}t^{2}(1-3t^{2})^{3})$.
If these conditions hold then $P\simeq A_{-27}$, i.e. $P$ is isomorphic to the quadratic twist of $A$ along $k(\omega)$. A sextic twist of $P$ satisfies the above conditions if and only if $j(C)=3t^{2}$ for some $t\in k^{\times}$.
Proof.
The proof is similar to that of Corollary 6.11 and is omitted. ∎
Remark 6.13.
Corollaries 6.11 and 6.12 show that a $j$-invariant $j\in Y({\rm\mathbb{Q}})={\rm\mathbb{P}}^{1}({\rm\mathbb{Q}})$ lifts to a ${\rm\mathbb{Q}}$-rational point under the morphism $\pi_{2}$ (respectively $\pi_{6}$) from §4.3 if and only if $j\in-{\rm\mathbb{Q}}^{\times 2}$ (respectively $j\in 3{\rm\mathbb{Q}}^{\times 2}$). We also note that the equivalence of $(1)$ and $(2)$ holds even when $j(C)=1$.
Corollary 6.14.
Suppose $k\neq k(\omega)$ and $P$ is geometrically simple. Then
  1. (1)
    ${\rm End}(P)$ is isomorphic to either ${\rm\mathbb{Z}}$, ${\rm\mathbb{Z}}[\sqrt{2}]$ or ${\rm\mathbb{Z}}[\sqrt{6}]$.
  2. (2)
    $P$ carries a principal polarization if and only if ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{2}]$, and in this case the unique principal polarization is $\frac{1}{2}\lambda(2-\sqrt{2})$.
Proof.
$(1)$ follows from Lemma 4.2. For $(2)$, if ${\rm End}(P)={\rm\mathbb{Z}}$, then ${\rm Hom}(P,A)={\rm\mathbb{Z}}\lambda$ so cannot contain a principal polarization. Otherwise, we use the fact that any abelian surface with ${\rm\mathbb{Z}}[\sqrt{2}]$-RM is principally polarized, whereas any $(1,2)$-polarized abelian surface $P$ with ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{6}]$ is not, by [19, Proposition 3.11]. If ${\rm End}(P)={\rm\mathbb{Z}}[\sqrt{2}]$, the unique principal polarization $\lambda^{\prime}\colon P\to A$ is $\frac{1}{2}\lambda(2-\sqrt{2})$; see [19, Proposition 2.1]. Equivalently, $\lambda^{\prime}=\lambda-\psi^{-1}$, in the notation of the proof of Proposition 6.1. ∎
6.4. Geometrically split Prym varieties
In this subsection, we assume that $k$ has characteristic zero. Let $C/k$ be a bielliptic Picard curve with Prym variety $P$. If any of the equivalent conditions of Lemma 6.2 is satisfied for $P_{\bar{k}}$ we call $C$ (or $P$) CM. Being CM only depends on the $j$-invariant (2.3) of $C$. Choose an embedding $\iota\colon\mathcal{O}\hookrightarrow{\rm End}(P_{\bar{k}})$ of the form $\iota_{x}$ for some $x\in\mathcal{S}$ as in §6.1. Let
(6.5) ${\rm End}_{\mathcal{O}}(P_{\bar{k}}):=\{f\in{\rm End}(P_{\bar{k}})\mid f\circ\iota(b)=\iota(b)\circ f\text{ for all }b\in\mathcal{O}\}.$
This subring does not depend on the choice of $x\in\mathcal{S}$. If $P$ is not CM, then ${\rm End}_{\mathcal{O}}(P_{\bar{k}})={\rm\mathbb{Z}}$; if $P$ is CM, then ${\rm End}_{\mathcal{O}}(P_{\bar{k}})$ is an order $R$ in an imaginary quadratic field, and we say that $P$ has CM by the order $R$.
Example 6.15.
The Prym variety $P$ of the special bielliptic Picard curve $C:y^{3}=x^{4}+1$ has CM by the order ${\rm\mathbb{Z}}[i]$, hence $j(C)=1$ is a CM $j$-invariant. (In this example alone, the symbol $i$ means $\sqrt{-1}$.) Indeed, over $\bar{{\rm\mathbb{Q}}}$, the curve $C$ has an automorphism $\beta(x,y)=(ix,y)$ commuting with the $\mu_{6}$-action on $C$. A signature calculation similar to Lemma 2.13 shows that $\beta$ induces an order $4$ automorphism $\beta_{*}$ on $P$ and has characteristic polynomial $X^{2}+1$. In the notation of §6.1, choose a pair $(\zeta,\varepsilon)\in\mathcal{S}$. By construction, the elements $r_{\zeta}$ and $s_{\varepsilon}$ commute with $\beta_{*}$, so ${\rm\mathbb{Z}}[\beta_{*}]={\rm\mathbb{Z}}[i]\subset{\rm End}_{\mathcal{O}}(P_{\bar{k}})$.
We now state some basic properties of the set of CM $j$-invariants, completely analogous to the elliptic curve setting, using the results of §4.3. In the notation of that subsection, a $j$-invariant $j\in Y({\rm\mathbb{C}})$ is CM if and only if it lifts to a CM point of $X$ under the map $\pi_{2}$.
Lemma 6.16.
Every CM $j$-invariant is algebraic over ${\rm\mathbb{Q}}$. Moreover, for every $d\in{\rm\mathbb{Z}}_{\geq 1}$, there are only finitely many CM $j$-invariants $j$ with $[{\rm\mathbb{Q}}(j):{\rm\mathbb{Q}}]\leq d$.
Proof.
Follows immediately from Proposition 4.7 and the corresponding facts concerning CM points Shimura curves [13, §2.4]. ∎
We conclude by determining all ${\rm\mathbb{Q}}$-rational CM $j$-invariants, equivalently all geometrically split $\operatorname{Prym}(C,\gamma)$ that can be defined over ${\rm\mathbb{Q}}$. To this end, we will use Elkies’ computations of all rational CM points on the full Atkin–Lehner quotient $X^{*}:=X/W\simeq{\rm\mathbb{P}}^{1}_{{\rm\mathbb{Q}}}$. Following Elkies [13, §3.1], let $t$ be the unique coordinate $X^{*}\xrightarrow{\sim}{\rm\mathbb{P}}^{1}_{t}$ with the property that $t=0,1,\infty$ corresponds to the unique CM point by the order ${\rm\mathbb{Z}}[\sqrt{-6}]$, ${\rm\mathbb{Z}}[i]$, ${\rm\mathbb{Z}}[\omega]$ respectively. Recall from Lemma 4.3 the isomorphism $j\colon Y\rightarrow{\rm\mathbb{P}}^{1}$. The isomorphism $Y\simeq(X/w_{3})^{\bar{w}}$ of Proposition 4.7 determines a quotient map $\Pi\colon Y\simeq(X/w_{3})^{\bar{w}}\rightarrow((X/w_{3})^{\bar{w}})/\langle\bar{w}\rangle=X^{*}$.
Lemma 6.17.
In the coordinates $j$ and $t$ of $Y$ and $X^{*}$ respectively, the map $\Pi\colon Y\rightarrow X^{*}$ is given by $j\mapsto\frac{(j+1)^{2}}{4j}$. A $k$-point $t\in k$ on $X^{*}$ lifts to $Y$ if and only if $t(t-1)$ is a square in $k$.
Proof.
Elkies [13, §3.1, end of p 17] has shown that the biquadratic extension of function fields ${\rm\mathbb{Q}}(X)/{\rm\mathbb{Q}}(X^{*})$ is given by adjoining $\sqrt{-t}$ and $\sqrt{3(t-1)}$ to ${\rm\mathbb{Q}}(X^{*})={\rm\mathbb{Q}}(t)$. The three intermediate quadratic subfields correspond to the three intermediate Atkin–Lehner quotients of $X$; analyzing the ramification shows that the morphism $X/w_{3}\rightarrow X^{*}$ can be realized as the double cover of $X^{*}={\rm\mathbb{P}}^{1}_{t}$ branched along the function $-3t(t-1)$, so $X/w_{3}$ has equation $s^{2}=-3t(t-1)$. Moreover, the involution $(t,s)\mapsto(t,-s)$ corresponds to the involution $\bar{w}$. Proposition 4.7 shows that $Y\rightarrow X^{*}$ is the quadratic twist of $X/w_{3}$ along ${\rm\mathbb{Q}}(\sqrt{-3})/{\rm\mathbb{Q}}$, so $Y\rightarrow{\rm\mathbb{P}}_{t}^{1}$ has equation $s^{2}=t(t-1)$. Rationally parametrizing this conic shows that $Y$ has a rational coordinate $r$ such that $s=(1-r^{2})/(4r)$, $t=(1+r)^{2}/(4r)$ and the involution $\bar{w}$ corresponds to $r\mapsto 1/r$.
We claim that after possibly replacing $r$ by $1/r$, which does not affect the expression $t=(1+r)^{2}/(4r)$, we have $j=r$. Analyzing the ramification of $\Pi$, we see that $Y$ has two CM points by ${\rm\mathbb{Z}}[\omega]$ (corresponding to $r=0,\infty$) and one CM point by ${\rm\mathbb{Z}}[i]$ (corresponding to $r=1$). The points $j=0,\infty$ correspond to decomposable $\mu_{3}$-abelian surfaces, described in the proof of Lemma 4.3, and have CM by ${\rm\mathbb{Z}}[\omega]$. Example 6.15 shows that $j=1$ corresponds to the CM point by ${\rm\mathbb{Z}}[i]$. So $j$ and $r$ (or $1/r$) agree on the three points $0,1,\infty$, so they must agree everywhere, proving the claim. ∎
Proposition 6.18.
Table 1 is a complete list of rational CM $j$-invariants and the discriminants of the corresponding quadratic orders. Consequently, a bielliptic Picard curve $C/{\rm\mathbb{Q}}$ has geometrically non-simple Prym variety if and only if $j(C)$ appears in this table.
Proof.
A point in $Y({\rm\mathbb{Q}})$ is CM if and only if its image under $\Pi\colon Y\rightarrow X^{*}$ is CM. It therefore suffices to find all rational CM points on $X^{*}$ and determine which ones lift to $Y({\rm\mathbb{Q}})$. Elkies [13, Table 1] has determined all rational CM points in his $t$-coordinate. Some of these were only conjecturally CM, but [14] confirms this table unconditionally. The proposition then follows from Lemma 6.17 and an elementary calculation. ∎
${\rm Disc}(R)$ $|{\rm Disc}(R)|$ $j$ or $1/j$
$-3$ $3$ 00
$-4$ $2^{2}$ $1$
$-24$ $2^{3}\cdot 3$ $-1$
$-84$ $2^{2}\cdot 3\cdot 7$ $-27$
$-120$ $2^{3}\cdot 3\cdot 5$ $27/125$
$-228$ $2^{2}\cdot 3\cdot 19$ $15625/729$
$-147$ $3\cdot 7^{2}$ $-48384/15625$
$-372$ $2^{2}\cdot 3\cdot 31$ $-1771561/421875$
$-408$ $2^{3}\cdot 3\cdot 17$ $-11390625/4913$
Table 1. Rational CM $j$-invariants and their discriminants
7. $6$-torsion points in the Prym variety
To prove Theorem 1.1, we must analyze the Galois modules $P[n]$ of bielliptic Picard Pryms. In this section we study $P[2]$, $P[3]$ and $P[6]$ explicitly and give various criteria for the existence of rational torsion points.
For the remainder of this section, let $(C,\gamma)=(C_{a,b},\gamma_{a,b})$ be a marked bielliptic Picard curve over a field $k$ (always assumed of characteristic $\neq 2,3$).
7.1. $2$-torsion
Recall from §2.5-2.6 that $[2]=\widehat{\lambda}\lambda$, giving rise to a short exact sequence
(7.1) $0\rightarrow P[\lambda]\rightarrow P[2]\rightarrow A[\widehat{\lambda}]\rightarrow 0$
The diagram from §2.5 shows that there is a canonical isomorphism $P[\lambda]\simeq E[2]$ and by bigonal duality we have $A[\widehat{\lambda}]\simeq\widehat{E}[2]$ as well. We conclude:
Lemma 7.1.
There is a short exact sequence of ${\rm Gal}_{k}$-modules
(7.2) $0\to E[2]\to P[2]\to\widehat{E}[2]\to 0$
The $k$-points of the outer terms in (7.2) are easy to determine.
Lemma 7.2.
Let $E=E_{a,b}$ and $\widehat{E}=\widehat{E}_{a,b}$ be as above. Then
  1. (1)
    $E[2](k)\neq 0$ if and only if $16(a^{2}-4b)$ is a cube in $k$, in which case $P[2](k)\neq 0$.
  2. (2)
    $\widehat{E}[2](k)\neq 0$ if and only if $b$ is a cube in $k$, in which case $A[2](k)\neq 0$.
Moreover if $k(\omega)\neq k$, then $|E[2](k)|\leq 2$ and $|\widehat{E}[2](k)|\leq 2$.
Proof.
This can be read off the models $E\colon y^{2}=x^{3}+16(a^{2}-4b)$ and $\widehat{E}\colon y^{2}=x^{3}+b$ of §2.7. ∎
[…]
Proposition 7.3.
Each divisor class in $P[2](\bar{k})$ is represented by a unique divisor of the form $R+\pi^{*}(T)-3\infty$, where $R\in C(\overline{k})$ is a ramification point of the map $\pi\colon C\rightarrow E$ and $T\in E(\overline{k})$ is such that $[2](T)=-\pi(R)$.
Proof.
We first show that every such divisor defines an element of $P[2](\bar{k})$, so let $x=R+\pi^{*}(T)-3\infty$ be such a divisor class. Recall that $\pi(\infty)=O_{E}$ denotes the origin of $E$. The condition $[2](T)=-\pi(R)$ translates to an equivalence $2T+\pi(R)\sim 3O_{E}$. Pulling this equivalence back along $\pi$ shows that $02x=$ 0, so $x\in J[2](\bar{k})$. Since both $R$ and $\pi^{*}(T)$ are fixed by $\tau$, $\tau(x)=x$ so $\tau(x)+x=2x=0$ so $x\in P[2](\bar{k})$. We now claim that two such divisors $R+\pi^{*}(T)-3\infty$ and $R^{\prime}+\pi^{*}(T^{\prime})-3\infty$ are linearly equivalent if and only if $R=R^{\prime}$ and $T=T^{\prime}$. Indeed, Proposition 2.15(1) shows that $R=R^{\prime}$, and then $T=T^{\prime}$ follows from the fact that $\pi^{*}\colon E\rightarrow J$ is injective.
Since there are $16$ divisors of this form and $P[2](\bar{k})$ has order $16$ too, every element of $P[2](\bar{k})$ has a unique representative of this form. ∎
This description of $P[2]$ is compatible with the sequence (7.2): the map $E[2]\rightarrow P[2]$ sends $T$ to $\pi^{*}(T)-2\infty$; the map $P[2]\rightarrow\widehat{E}[2]$ sends $R+\pi^{*}(T)-3\infty$ with $R=(0,t)\in C(\bar{k})$ to $(-t,0)\in\widehat{E}[2](\bar{k})$, using the coordinates of the equations in §2.7.
Theorem 7.4.
An element $(-t,0)\in\widehat{E}[2](k)$ with $t\in k$ lifts to $P[2](k)$ under (7.2) if and only if $g_{a,t}(z):=z^{4}-6tz^{2}+4az-3t^{2}$ has a $k$-rational root. Consequently, $(P[2]\setminus P[\lambda])(k)\neq\varnothing$ if and only if there exists $t\in k$ such that $t^{3}=b$ and such that $g_{a,t}$ has a $k$-rational root.
Proof.
By the above proposition, the point lifts to $P[2](k)$ if and only if the corresponding point $(4t,4a)\in E(k)$ is divisible by $2$ in $E(k)$. Using the $2$-descent map $E(k)/2E(k)\hookrightarrow\operatorname{H}^{1}(k,E[2])$ and its interpretation via binary quartic forms, this is equivalent to $z^{4}-24tz^{2}+32az-48t^{2}$ having a root in $k$ [2, §2, (11)]. Changing $z$ by $2z$ and dividing by $16$ results in the polynomial $g_{a,t}$. ∎
Corollary 7.5.
$(P[2]\setminus P[\lambda])(k)\neq\varnothing$ if and only if $(a,b)=((4s+3)(4s^{2}-3)d^{3},(4s+3)^{3}d^{6})$ for some $s,d\in k$.
Proof.
Theorem 7.4 shows that $P[2]\setminus P[\lambda]$ has a $k$-point if and only if there exists a $t\in k$ with $t^{3}=b$ such that $g_{a,t}(z)=z^{4}-6tz^{2}+4az-3t^{2}$ has a $k$-rational root. A calculation shows that if $(a,b)$ is of the above form for some $s,d\in k$ then $t=(4s+3)$ is such that $g_{a,t}(z)$ has the root $z=-4s-3$ in $k$. Conversely, suppose that $t\in k$ is a cube root of $b$ such that $g_{a,t}(z)$ has a root in $k$. If $a=0$, a calculation shows that we may take $s=\pm\sqrt{3}/2$, so we may assume that $a\neq 0$. Note that $\lambda g_{a,t}(z)=g_{\lambda^{3}a,\lambda^{2}t}(\lambda z)$ for all $\lambda\in k$. Choosing $\lambda=t/a$ and setting $v=t^{3}/a^{2}$, it follows that $g_{v,v}(z)$ has a $k$-rational root. Since $t\neq 0$, this root must be nonzero. The locus of $(v,z)$ with $g_{v,v}(z)=0$ and $z\neq 0$ is isomorphic to an open subset of a smooth conic. After rationally parametrizing this conic we find that $v$ must be of the form $(4s+3)/(4s^{2}-3)^{2}$ for some $s\in k$. It follows that $(a,b)$ is of the above form with $d=a/t$. ∎
In the following we assume for simplicity that $k(\omega)\neq k$, so that $E[2](k)\neq E[2]$. For abstract groups $G$ and $G^{\prime}$, we sometimes write $G\subset G^{\prime}$ as shorthand for “there exists an embedding $G\hookrightarrow G^{\prime}$”.
Proposition 7.6.
Suppose $k(\omega)\neq k$. Then the following conditions are equivalent:
  1. (1)
    $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\subset P(k)$;
  2. (2)
    $E[2](k)\neq 0$ and there exists a branch point of $C\rightarrow E$ different from $O_{E}$ lying in $2E(k)$;
  3. (3)
    $(a,b)=((16w^{6}+40w^{3}-2)d^{3},\left(8w^{3}+1\right)^{3}d^{6})$ for some pair $w,d\in k$;
  4. (4)
    $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\subset A(k)$.
Proof.
That $(1)$ and $(2)$ are equivalent follows from (7.2), Lemma 7.2, the fact that $k(\omega)\neq k$ and Proposition 7.3. Similarly, Proposition 7.6 shows that $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\subset P_{a,b}(k)$ if and only if $(a,b)=((4s+3)(4s^{2}-3)d^{3},(4s+3)^{3}d^{6})$ for some $s,d\in k$ and $16(a^{2}-4b)=16d^{6}(2s-3)(2s+1)^{3}(4s+3)^{2}$ is a cube. The latter is equivalent to $(3-2s)/4(4s+3)$ being a cube $w^{3}$ in $k$. Solving for $s$, plugging in back to $a,b$ and absorbing common factors in $d$ proves the equivalence between $(1)$ and $(3)$.
A calculation shows that if $(a,b)$ is of the form $(3)$ for some $w,d\in k$, then $(8a,16(a^{2}-4b))$ is of the form $(3)$ with $w^{\prime}=-1/2w$ and $d^{\prime}=-(2t)^{2}d$. Therefore $(3)\Leftrightarrow(4)$ by bigonal duality from §2.7. ∎
[…]
7.2. $\sqrt{-3}$-torsion
We use the $\mu_{3}$-action on $P$ to explicitly analyze a ${\rm Gal}_{k}$-submodule of $P[3]$. Let $\mathfrak{p}$ be the ideal $(1-\omega)\subset{\rm\mathbb{Z}}[\omega]$, viewed as a subgroup of ${\rm End}(P_{k^{\operatorname{sep}}})$. Since $\mathfrak{p}$ is ${\rm Gal}_{k}$-stable, so is the subgroup $A[\mathfrak{p}]$ of points fixed by $\omega$. We have $\mathfrak{p}=(\sqrt{-3})$, so $P[\mathfrak{p}]$ has order $9$, and is the kernel of an isogeny $P\to B$ over $k$. In fact, the quotient $B=P/P[\mathfrak{p}]$ is isomorphic to the sextic twist $P_{-27}$, which is also the quadratic twist of $P$ by the extension $k(\omega)/k$ [41, Rem 2.8]. We therefore have an exact sequence
(7.3) $0\to P[\mathfrak{p}]\to P[3]\to P_{-27}[\mathfrak{p}]\to 0$
Of course, $A=P^{\vee}$ also has a $\mu_{3}$-action, so we can similarly define $A[\mathfrak{p}]\subset A[3]$, a subgroup of order $9$ that sits in an analogous short exact sequence.
Lemma 7.7.
$A[\mathfrak{p}]\simeq P[\mathfrak{p}]$ as ${\rm Gal}_{k}$-modules.
Proof.
The restriction of the polarization $\lambda$ induces such an isomorphism. ∎
Let $f(x):=x^{4}+ax^{2}+b\in k[x]$ and let ={±$\mathcal{R}=\{\pm\alpha_{1},\pm\alpha_{2}\}$ be the $4$ roots of $f$ in $k^{\operatorname{sep}}$.
Lemma 7.8.
The Galois module $P[{\rm\mathfrak{p}}]$ is generated by the classes $D_{1}:=(\alpha_{1},0)-(-\alpha_{1},0)$ and $D_{2}:=(\alpha_{2},0)-(-\alpha_{2},0)$.
Proof.
These classes live in $P$, are non-trivial, and are annihilated by $1-\omega$ so lie in $P[{\rm\mathfrak{p}}]$. They are linearly independent over ${\rm\mathbb{F}}_{3}$ since otherwise we would have $D_{2}\sim\pm D_{1}$, which is impossible since $C$ is not hyperelliptic. ∎
Let $D_{4}\subset{\rm Sym}(\mathcal{R})$ the subset of permutations of $\mathcal{R}$ such that $(-\alpha)^{\sigma}=-\alpha^{\sigma}$ for all $\alpha\in\mathcal{R}$ and $\sigma\in{\rm Sym}(\mathcal{R})$. This is a dihedral group of order $8$, generated by its four reflections $\{\tau_{1},\tau_{2},\hat{\tau}_{1},\hat{\tau}_{2}\}$, where $\tau_{i}$ maps $\alpha_{i}$ to $-\alpha_{i}$ and fixes the other roots, $\hat{\tau}_{1}$ swaps $\alpha_{1}\leftrightarrow\alpha_{2}$ and $-\alpha_{1}\leftrightarrow-\alpha_{2}$, and $\hat{\tau}_{2}$ swaps $\alpha_{1}\leftrightarrow-\alpha_{2}$ and $-\alpha_{1}\leftrightarrow\alpha_{2}$. (The reader is invited to picture $D_{4}$ acting on a square with labels $\{\alpha_{1},\alpha_{2},-\alpha_{1},-\alpha_{2}\}$.) Let $V$ be the ${\rm\mathbb{F}}_{3}$-vector space with basis $\{v_{\alpha}\mid\alpha\in\mathcal{R}\}$, modulo the relations $v_{-\alpha}=-v_{\alpha}$. The $D_{4}$-action on $\mathcal{R}$ induces a linear $D_{4}$-action on $V$. After choosing the basis $\{v_{\alpha_{1}},v_{\alpha_{2}}\}$, this induces a representation $\varphi\colon D_{4}\rightarrow{\rm GL}_{2}({\rm\mathbb{F}}_{3})$, isomorphic to the mod $3$ reduction of the reflection representation of $D_{4}$. The ${\rm Gal}_{k}$-action on $\mathcal{R}$ determines a homomorphism $\chi\colon{\rm Gal}_{k}\rightarrow{\rm GL}_{2}({\rm\mathbb{F}}_{3})$. Lemma 7.8 shows:
Proposition 7.9.
[…]
Corollary 7.10.
We have $P[{\rm\mathfrak{p}}](k)\simeq({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}$ if and only if $f(x)$ splits completely in $k[x]$.
Proof.
Use Proposition 7.9 and the fact that $\varphi$ is injective. ∎
Corollary 7.11.
The ${\rm\mathbb{F}}_{3}[{\rm Gal}_{k}]$-module $P[{\rm\mathfrak{p}}]$ is semisimple.
Proof.
Use Proposition 7.9 and the fact that $3$ is coprime to the order of $D_{4}$. ∎
Let $\hat{f}:=x^{4}+8ax+16(a^{2}-4b)$.
Corollary 7.12.
$P[\mathfrak{p}](k)\neq 0$ if and only if either $f$ or $\hat{f}$ has a root in $k$.
Proof.
Let $G$ be the image of $\chi\colon{\rm Gal}_{k}\rightarrow D_{4}$. Then $P[\mathfrak{p}](k)\neq 0$ if and only if $G$ fixes some nonzero element of $V={\rm\mathbb{F}}_{3}^{2}$. A group theory calculation shows that only reflections have fixed points, so this is equivalent to $G\subset\langle\tau_{i}\rangle$ or $G\subset\langle\hat{\tau}_{i}\rangle$ for some $i=1,2$. We have $G\subset\langle\tau_{i}\rangle$ for some $i$ if and only if $f$ has a $k$-rational root. By Lemma 2.21, $G\subset\langle\hat{\tau}_{i}\rangle=\text{Stab}_{D_{4}}(\alpha_{1}+\alpha_{2})$ for some $i$ if and only if $\hat{f}$ has a $k$-rational root. ∎
Proposition 7.13.
${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\subset P[\mathfrak{p}](k)$ if and only if either
  1. (1)
    $P=P_{-(c+1)d^{2},cd^{4}}$ for some $c,d\in k$ or
  2. (2)
    $P=P_{-8(c+1)d^{2},16(c-1)^{2}d^{4}}$ for some $c,d\in k$.
Proof.
The polynomial $f(x)$ has a $k$-rational root if and only if $f(x)=(x^{2}-t^{2})(x^{2}-c)$ for some $c,t\in k$, if and only if $P\simeq P_{-c-t^{2},ct^{2}}$ for some $c,t\in k$ if and only if $P$ is as in $(1)$. Similarly $\widehat{f}$ has a $k$-root if and only if $P$ is as in $(2)$. The proposition then follows from Corollary 7.12. ∎
Proposition 7.14.
$({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}\subset P[\mathfrak{p}](k)$ if and only if $P=P_{-(c^{2}+1)d^{2},c^{2}d^{4}}$ for some $c,d\in k$.
Proof.
By Corollary 7.10, this happens if and only if $f(x)$ splits completely, in which case we have $f(x)=(x^{2}-t^{2})(x^{2}-c^{2})$ for some $c,t\in k$. Up to cubic twist we may take $t=1$, hence we arrive at the form $P_{-(c^{2}+1)d^{2},c^{2}d^{4}}$, for some $c,d\in k$. ∎
Question 1.
Do there exist bielliptic Picard curves $C/{\rm\mathbb{Q}}$ such that $P[3]({\rm\mathbb{Q}})\neq 0$ but $P[{\rm\mathfrak{p}}]({\rm\mathbb{Q}})=0$?
7.3. $6$-torsion
We combine our results on $P[2]$ and $P[3]$ to produce examples of points of order $6$ in $P(k)$. Here we assume $k={\rm\mathbb{Q}}$ for simplicity.
Proposition 7.15.
${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}\subset P[2\mathfrak{p}]({\rm\mathbb{Q}})$ if and only if one of the following holds
  1. (1)
    $P=P_{2^{4}(c+1)(c-1)^{2},2^{8}c(c-1)^{4}}$ for some $c\in{\rm\mathbb{Q}}$; or
  2. (2)
    $P=P_{8c(1-c),16c^{2}(1+c)^{2}}$ for some $c\in{\rm\mathbb{Q}}$.
  3. (3)
    $P=P_{-(c+1)c^{4},16(1-c)^{2}c^{8}}$ where $c=\frac{(3t-2)(5t-2)^{3}}{t(7t-4)^{3}}$ for some $t\in{\rm\mathbb{Q}}$.
  4. (4)
    $P=P_{\frac{1}{4}(1-v)^{3}(3v+1)^{3}(3v^{4}+6v^{2}-1),v^{6}(v-1)^{6}(3v+1)^{6}}$ for some $v\in{\rm\mathbb{Q}}$.
Proof.
Each family above is the fiber product of one of the two families of rational $2$-torsion in $P$ (Lemma 7.2 and Corollary 7.5) with one of the two families in Proposition 7.13. We omit the details, as it is tedious but straightforward algebra, in a spirit similar to Proposition 7.6. ∎
Proposition 7.16.
${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/6{\rm\mathbb{Z}}\subset P[2{\rm\mathfrak{p}}]({\rm\mathbb{Q}})$ if and only if there exists $c\in{\rm\mathbb{Q}}\setminus\{0,\pm 1\}$ such that $a=-16(c^{2}+1)(c^{2}-1)^{2}$ and $b=2^{8}c^{2}(c^{2}-1)^{4}$.
Proof.
This follows by combining Proposition 7.14 with both Lemma 7.2 and Corollary 7.5. In the first case, we are forced to set $d=4(c^{2}-1)$ which leads to the formulas for $a$ and $b$ above. In the second case, we must first set $d=c$ in order for $b$ to be a cube, with $b=t^{3}$ where $t=c^{2}$. Then we must see when the polynomial
$g_{a,t}(z)=g_{c}(z)=z^{4}-6c^{2}z^{2}-4c^{4}z-4c^{2}z-3c^{4}$
has a root. The plane curve $\{(c,z)\colon g_{c}(z)=0\}$ is irreducible of geometric genus $1$, and with the help of Magma [7] we find that it is birational to an elliptic curve with Mordell-Weil group of order $8$. None of these rational points correspond to smooth bielliptic Picard curves, so this second case gives no new examples. ∎
Simple algebra shows that for $(a,b)$ as in Proposition 7.16, the curve $C_{a,b}$ has affine model
$2(c^{2}-1)^{2}y^{3}=(x^{2}-1)(x^{2}-c^{2}),$
recovering the equation given in the introduction.
Remark 7.17.
Combining Proposition 7.16 and Remark 2.12, we are led to the family of genus two curves
$C_{t}\colon(t^{2}+1)y^{2}=(2x^{2}+2x-1)((t^{2}-1)^{2}x^{4}+2(t^{2}-1)^{2}x^{3}+4t^{2}x-t^{2}).$
For all but finitely many $t\in{\rm\mathbb{Q}}$, the Jacobian $J=J_{t}=\operatorname{Jac}(C_{t})$ satisfies ${\rm End}^{0}(J_{\overline{{\rm\mathbb{Q}}}})\simeq B$. For all integers $1\leq t\leq 10000$ we compute $J_{t}({\rm\mathbb{Q}})_{\mathrm{tors}}\simeq{\rm\mathbb{Z}}/2{\rm\mathbb{Z}}\times({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}$ in Magma [7]. We suspect this holds for all (but finitely many) $t\in{\rm\mathbb{Q}}$, since the isogeny of Remark 2.12 should have degree $2$.
7.4. Infinite families
Table 2 summarizes many of the computations of Section §7. For each group $G$ in Theorem 1.1, the table gives an infinite family of curves $C_{a,b}$ such that $G\hookrightarrow P({\rm\mathbb{Q}})_{\mathrm{tors}}$. Since the $j$-invariant of every family is not constant, each family contains infinitely many distinct $\bar{{\rm\mathbb{Q}}}$-isomorphism classes, and all but finitely many of these are geometrically simple by Lemma 6.16. For certain groups $G$ there exist multiple such families, but we only write one. In particular, we do not claim that every Prym $P$ such that $G\hookrightarrow P({\rm\mathbb{Q}})_{\mathrm{tors}}$ appears in this table.
$G$ $a$ $b$ $j$
$\{1\}$ $a$ $b$ $\frac{4b-a^{2}}{4b}$
${\rm\mathbb{Z}}/2{\rm\mathbb{Z}}$ $2s$ $s^{2}-t^{3}$ $\frac{t^{3}}{t^{3}-s^{2}}$
${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$ $-(c+1)d^{2}$ $cd^{4}$ $\frac{-(c-1)^{2}}{4c}$
$({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}$ $(16w^{6}+40w^{3}-2)d^{3}$ $\left(8w^{3}+1\right)^{3}d^{6}$ $-\left(\frac{4w\left(w^{3}-1\right)}{\left(8w^{3}+1\right)}\right)^{3}$
${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$ $8c(1-c)$ $16c^{2}(1+c)^{2}$ $\frac{4c}{(c+1)^{2}}$
$({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}$ $-(c^{2}+1)d^{2}$ $c^{2}d^{4}$ $\frac{(c^{2}-1)^{2}}{-4c^{2}}$
${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$ $-16(c^{2}+1)(c^{2}-1)^{2}$ $2^{8}c^{2}(c^{2}-1)^{4}$ $\frac{(c^{2}-1)^{2}}{-4c^{2}}$
Table 2. Infinite families of Pryms $P_{a,b}$ with $G\hookrightarrow P_{a,b}({\rm\mathbb{Q}})_{\mathrm{tors}}$.
7.5. $2$-torsion in $A$ using bitangents
Even though it is possible to use the description of $P[2]$ from §7.3 and bigonal duality to describe $A[2]$, we give a more intrinsic description of the ${\rm Gal}_{k}$-module using the bitangents of the quartic curve $C$. This will be used in §8.3 to rule out rational points of order $4$ in $P$.
We first recall the connection between the bitangents on a smooth plane quartic curve $C$ over a general field $k$ (of characteristic not $2$ nor $3$) and $2$-torsion points in its Jacobian; see [11, §6] for more details. Let $X\subset{\rm\mathbb{P}}^{2}_{k}$ be a smooth plane quartic curve. A bitangent of $X$ is a line $\ell\subset{\rm\mathbb{P}}^{2}_{k}$ which intersects $X$ with even multiplicity at every point, i.e. $\ell\cap X=2D$ for some effective divisor $D$ of $X$ of degree $2$. Since $X$ is canonically embedded, $2D\sim K_{X}$ and so $D$ is a theta characteristic. This sets up a bijection between the $28$ bitangents of $X_{\bar{k}}$ and the odd theta characteristics of $X_{\bar{k}}$. If $D,E$ are two odd theta characteristics of $X_{\bar{k}}$, then $D-E\in\operatorname{Jac}_{X}[2](\bar{k})$. Every element of $\operatorname{Jac}_{X}[2](\bar{k})$ is a sum of points of this form.
We now specialize to our setting where $(C,\gamma)$ is a bielliptic Picard curve over $k$. Note that the line at infinity is a bitangent with theta characteristic $2\infty$.
Lemma 7.18.
The action of $\langle\tau\rangle$ on the $28$ bitangents of $C_{\bar{k}}$ has four fixed points (normal-(((including the line at infinity)normal-))) and $12$ orbits of size $2$.
Proof.
Using the explicit equation (2.1), $\tau(x,y)=(-x,y)$. Therefore $\tau$ is fixes a bitangent line (different from the line at infinity) if and only if it is of the form $y=c$. Such an equation defines a bitangent line if and only if $x^{4}+ax^{2}+b-c$ is the square of a polynomial, which happens if and only if $c^{3}=b-a^{2}/4$, which has $3$ solutions in $\bar{k}$. ∎
Write $p\colon J\rightarrow A$ for the projection map, the dual of $P\hookrightarrow J$. Given any bitangent $\ell$ of $C_{\bar{k}}$ with theta characteristic $D$, write $x_{\ell}=D-2\infty$ for the associated $2$-torsion point in $J(\bar{k})$.
Proposition 7.19.
If $\ell$ is a bitangent line fixed by $\tau$, then $p(x_{\ell})=0$. If $\ell$ and $\ell^{\prime}$ are two distinct bitangent lines of $C$ not fixed by $\tau$, then $p(x_{\ell})=p(x_{\ell^{\prime}})$ if and only if $\ell^{\prime}=\tau(\ell)$. The map $\ell\mapsto p(x_{\ell})$ induces a bijection between the $\langle\tau\rangle$-orbits of bitangents of $C_{\bar{k}}$ of size $2$ and $A[2](\bar{k})\setminus A[\widehat{\lambda}](\bar{k})$.
Proof.
We may assume that $k=\bar{k}$. If $\ell$ is fixed by $\tau$, then $\ell\cap C=2\pi^{-1}(R)$ for some $R\in E[2]$. Since the morphism $p$ has kernel $\pi^{*}(E)$, $p(x_{\ell})=0$. We claim that if $\ell$ is not fixed by $\tau$ then $p(x_{\ell})\not\in A[\widehat{\lambda}]$. Since $J\simeq(P\times E)/(P\cap\pi^{*}(E))$, this is equivalent to saying that $x_{\ell}\notin P[2]$. This follows from the assumption that $D\not\sim\tau(D)$ and since $\tau$ acts as $-1$ on $P$. Since $A[2]\setminus A[\widehat{\lambda}]$ and the set of $\langle\tau\rangle$-orbits of bitangents of size $2$ both have size $12$, the remainder of the lemma follows from Proposition 2.16. ∎
8. Classifying rational torsion in Prym varieties
In Section 7 we showed that the finite groups mentioned in Theorem 1.1 all arise as subgroups of Pryms surfaces, and in fact infinitely often. In this section we finish the proof of the theorem by showing that all other finite abelian groups do not arise.
Let $C=C_{a,b}$ be a bielliptic Picard curve over ${\rm\mathbb{Q}}$, with double cover $\pi\colon C\to E$, Prym $P$, and dual Prym $A=P^{\vee}$, as usual.
8.1. Eliminating $\ell$-torsion for $\ell\geq 5$
We first show that $P[\ell]({\rm\mathbb{Q}})=0$ for all primes $\ell\geq 5$. When $P$ is geometrically simple, this follows from the more general result [28, Theorem 1.1]. We briefly explain how the techniques of that paper can be adapted to handle the geometrically split case as well and in the process give a simplified version of the proof (which is possible due to the specific nature of our family).
Fix a prime $\ell\geq 5$. Theorem 6.5 shows that there exists an embedding $\iota\colon\mathcal{O}\rightarrow{\rm End}(P_{\bar{{\rm\mathbb{Q}}}})$ whose image is ${\rm Gal}_{{\rm\mathbb{Q}}}$-stable; we will fix such an embedding in what follows. Let $\mathcal{O}_{\ell}:=\mathcal{O}\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{F}}_{\ell}$. Since $\mathcal{O}$ has discriminant $6$ and $\mathcal{O}$ is maximal at $\ell$, there is an isomorphism $\mathcal{O}_{\ell}\simeq\operatorname{Mat}_{2}({\rm\mathbb{F}}_{\ell})$. Let $M:=P[\ell](\bar{{\rm\mathbb{Q}}})$, a free left $\mathcal{O}_{\ell}$-module of rank $1$. Both $\mathcal{O}_{\ell}$ and $M$ are (right) ${\rm Gal}_{{\rm\mathbb{Q}}}$-modules, and the action of $\mathcal{O}_{\ell}$ on $M$ is ${\rm Gal}_{{\rm\mathbb{Q}}}$-equivariant.
Proposition 8.1.
If $P[\ell]({\rm\mathbb{Q}})\neq 0$, then $\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{Q}}(\omega)})$.
Proof.
The proof of [28, Theorem 6.0.1] carries over essentially without change; we briefly sketch the details. Suppose that $m\in M^{{\rm Gal}_{{\rm\mathbb{Q}}}}$ is nonzero. An argument identical to [28, Lemma 6.2.3] shows that $\mathcal{O}_{\ell}\cdot m\subset M$ has order $\ell^{2}$. Let $S:={\rm\mathbb{Z}}[\omega]\subset\mathcal{O}$ and $S_{\ell}:=S\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{F}}_{\ell}$. Then $S$ is a ${\rm Gal}_{{\rm\mathbb{Q}}}$-stable subring of $\mathcal{O}$, and the induced action of ${\rm Gal}_{{\rm\mathbb{Q}}}$ on $S$ factors through ${\rm Gal}({\rm\mathbb{Q}}(\omega)/{\rm\mathbb{Q}})$. Since $S_{\ell}$ has no ${\rm Gal}_{{\rm\mathbb{Q}}}$-stable proper nonzero ideals, the map $S_{\ell}\rightarrow\mathcal{O}\cdot m$ sending $x\mapsto x\cdot m$ is injective. By cardinality reasons, it is also surjective, so $S_{\ell}\cdot x=\mathcal{O}_{\ell}\cdot x$. The (purely linear-algebraic) [28, Lemma 6.2.6] then shows that ${\rm Gal}_{{\rm\mathbb{Q}}(\omega)}$ acts trivially on $\mathcal{O}_{\ell}$. Since $\ell\geq 5$, by [28, Lemma 3.5.7] this implies that ${\rm Gal}_{{\rm\mathbb{Q}}(\omega)}$ acts trivially on $\mathcal{O}$. In other words, $\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{Q}}(\omega)})$, as desired. ∎
The quaternionic multiplication places strong restrictions on the reduction type of $P$ at a prime $p$. First of all, $P$ has potentially good reduction at $p$, and acquires good reduction over a totally ramified extension $K$ of ${\rm\mathbb{Q}}_{p}$ [28, Lemma 4.1.2]. The special fiber of the Neron model of $P_{K}$ is an abelian surface over ${\rm\mathbb{F}}_{p}$; we will denote this abelian surface by $P_{{\rm\mathbb{F}}_{p}}$ and (by slight abuse of language) call it the reduction of $P$ mod $p$. (Its isomorphism class might depend on the choice of $K$.) The reduction $P_{{\rm\mathbb{F}}_{p}}$ is geometrically isogenous to the square of an elliptic curve. If $p$ divides the discriminant of $\mathcal{O}$ (that is, if $p\mid 6$), then $P_{{\rm\mathbb{F}}_{p}}$ is supersingular [24, §2]. Additionally, the prime-to-$p$ torsion subgroup of $P({\rm\mathbb{Q}})$ embeds in $P_{{\rm\mathbb{F}}_{p}}({\rm\mathbb{F}}_{p})$ for every prime $p$ by formal group considerations. We use these remarks to prove the following:
Proposition 8.2.
Let $\ell\geq 5$ be a prime. Then $P[\ell]({\rm\mathbb{Q}})=0$.
Proof.
Suppose instead that $P[\ell]({\rm\mathbb{Q}})\neq 0$. Then $\ell$ divides the order of $P_{{\rm\mathbb{F}}_{3}}({\rm\mathbb{F}}_{3})$. Moreover, $P_{{\rm\mathbb{F}}_{3}}$ is supersingular. Proposition 8.1 shows that $\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{Q}}(\omega)})$, so the base change $P_{{\rm\mathbb{Q}}(\omega)}$ has quaternionic multiplication over ${\rm\mathbb{Q}}(\omega)$. Since $3$ is ramified in ${\rm\mathbb{Q}}(\omega)$, we must also have $\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{F}}_{3}})$. By Honda-Tate theory (conveniently recorded in the results of the following LMFDB [31] search), we see that the only option is $\ell=7$.
We conclude by excluding the case $\ell=7$. Let $L/{\rm\mathbb{Q}}$ be the endomorphism field, namely the smallest field extension with the property that ${\rm End}(P_{L})={\rm End}(P_{\bar{{\rm\mathbb{Q}}}})$. Since ${\rm Gal}({\rm\mathbb{Q}}(\omega)/{\rm\mathbb{Q}})$ acts nontrivially on the subring ${\rm\mathbb{Z}}[\omega]\subset\mathcal{O}\subset{\rm End}(P_{{\rm\mathbb{Q}}(\omega)})$, we have ${\rm\mathbb{Q}}(\omega)\subset L$. A result of Silverberg [43, Theorem 4.2] shows that $L$ is unramified at all places of good reduction of $P$. Since $3$ is ramified in ${\rm\mathbb{Q}}(\omega)\subset L$, it follows that $P$ has bad reduction at $3$. By [28, Proposition 4.1.3(b)] and Lemma 6.9, the reduction of $P$ at $3$ is totally additive; in other words the identity component of the special fibre of the Neron model of $P$ over ${\rm\mathbb{Z}}_{3}$ is unipotent. A result of Lorenzini [32, Corollary 3.25] then shows that $7\nmid|P({\rm\mathbb{Q}}_{3})_{\mathrm{tors}}|$, a contradiction. ∎
Remark 8.3.
The elliptic curve $X\colon y^{2}=x^{3}+48(\omega+5)$ has a ${\rm\mathbb{Q}}(\omega)$-rational point of order $7$. It follows that the decomposable $\mu_{3}$-abelian surface $X\times X$ has (two independent) ${\rm\mathbb{Q}}(\omega)$-rational points of order $7$. This shows that the argument above is sharp in a certain sense.
8.2. Eliminating small groups
We rule out certain small groups of order $2^{i}3^{j}$ from appearing as subgroups of $P({\rm\mathbb{Q}})$, using arguments that are specific to our family.
Proposition 8.4.
${\rm\mathbb{Z}}/9{\rm\mathbb{Z}}\not\subset P({\rm\mathbb{Q}})$ and $P({\rm\mathbb{Q}})[3]\subset({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}$.
Proof.
Using the notation and remarks made before Proposition 8.2, the prime-to-$2$ torsion subgroup of $P({\rm\mathbb{Q}})$ injects into $P_{{\rm\mathbb{F}}_{2}}({\rm\mathbb{F}}_{2})$ and $P_{{\rm\mathbb{F}}_{2}}$ is a supersingular abelian surface. Consulting the LMFDB [31], we see that the $3$-part of $P_{{\rm\mathbb{F}}_{2}}({\rm\mathbb{F}}_{2})$ is of order at most $9$, and there is a unique isogeny class of supersingular abelian surfaces $B/{\rm\mathbb{F}}_{2}$ with $9\mid|B({\rm\mathbb{F}}_{2})|$. This isogeny class is the square of a supersingular elliptic curve $E$ over ${\rm\mathbb{F}}_{2}$ and $E({\rm\mathbb{F}}_{2})\simeq{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$. By [28, Lemma 7.2.1] (and its proof) and the fact that ${\rm\mathbb{Z}}[\sqrt{-2}]$ is a PID, every member of this isogeny class is in fact isomorphic to $E^{2}$, so $B({\rm\mathbb{F}}_{2})\simeq({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}$. We conclude that the $3$-part of $P_{{\rm\mathbb{F}}_{2}}({\rm\mathbb{F}}_{2})$, hence also the $3$-part of $P({\rm\mathbb{Q}})_{\mathrm{tors}}$, is a subgroup of $({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}$. ∎
Proposition 8.5.
$({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\times{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\not\subset P({\rm\mathbb{Q}})$.
Proof.
Suppose that $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\times({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})\subset P({\rm\mathbb{Q}})$. By Proposition 7.6, we have $j=-\left(\frac{4w\left(w^{3}-1\right)}{\left(8w^{3}+1\right)}\right)^{3}$ for some $w\in{\rm\mathbb{Q}}$. Since ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\subset P({\rm\mathbb{Q}})$, the exact sequence (7.3) shows that $P[\mathfrak{p}]({\rm\mathbb{Q}})\neq 0$ or $P_{-27}[\mathfrak{p}]({\rm\mathbb{Q}})\neq 0$. Proposition 7.13 then shows one of $j$ or $1/j$ is of the form $-4t/(t-1)^{2}$ for some $t\in{\rm\mathbb{Q}}$. Replacing $j$ by $1/j$ if necessary, using Proposition 7.6, Lemma 7.7 and bigonal duality, we may assume $j=-4t/(t-1)^{2}$ for some $t\in{\rm\mathbb{Q}}$. This is equivalent to
(8.1) $(4j-2)^{2}-4=\frac{1024w^{3}\left(w^{3}-1\right)^{3}\left(8w^{6}+20w^{3}-1\right)^{2}}{\left(8w^{3}+1\right)^{6}}$
being a square in ${\rm\mathbb{Q}}$. Since $w=0,1,-1/2$ leads to $j=0,\infty$, we may assume $w$ is distinct from these values. Then this expression is a square if and only if $z^{2}=w(w^{3}-1)$ for some $z\in{\rm\mathbb{Q}}$. This equation defines the affine part $C^{\circ}$ of a genus $1$ curve $C/{\rm\mathbb{Q}}$ which is a double cover of ${\rm\mathbb{P}}^{1}$. The point $(w,z)=(0,0)$ endows $C$ with the structure of an elliptic curve with Weierstrass model $y^{2}=x^{3}+1$. (This can be seen using the invariant theory of binary quartics [2, §2.1].) This elliptic curve has Mordell-Weil group ${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$, so $C^{\circ}({\rm\mathbb{Q}})=\{(0,0),(-2,0),(1,\pm 3)\}$. We conclude that $w\in\{0,1,-2\}$, but we already observed these all correspond to $j=0,\infty$. ∎
8.3. Eliminating points of order $4$
We rule out the existence of order $4$ points in $P({\rm\mathbb{Q}})$ using the description of $A[2]$ via bitangents from §7.5.
Proposition 8.6.
${\rm\mathbb{Z}}/4{\rm\mathbb{Z}}\not\subset P({\rm\mathbb{Q}})$.
Proof.
For the sake of contradiction, let $x\in P({\rm\mathbb{Q}})$ be a point of order $4$. Let $A=P^{\vee}$ be the dual Prym. We have $\widehat{\lambda}\circ\lambda=[2]$, so either $\lambda(x)\in A[2]$ or $y=\lambda(x)\in A({\rm\mathbb{Q}})$ has order $4$ and $\widehat{\lambda}(y)\in P[2]$. After possibly replacing $P$ by $A$ and applying bigonal duality (Proposition 2.19), we may therefore assume that $\lambda(x)\in A[2]({\rm\mathbb{Q}})$. Since $x$ has order $4$, we have $\lambda(x)\not\in A[\widehat{\lambda}]$. Using Proposition 7.19 and its notation, there exists a unique $\tau$-orbit of odd theta characteristics $\{D,D^{\prime}\}$ such that $p(D-2\infty)=\lambda(x)$. This orbit is defined over ${\rm\mathbb{Q}}$ but $D$ and $D^{\prime}$ may only be defined over a quadratic extension.
[…]
(8.2) $x-(D-2\infty)\sim\pi^{*}(y-O_{E})$
for some unique $y\in E(\bar{{\rm\mathbb{Q}}})$. The left hand side has order $4$, so $y$ has order $4$ too.
We claim that the subgroup $\{0,y,2y,3y\}$ generated by $y$ is ${\rm Gal}_{{\rm\mathbb{Q}}}$-stable. Indeed, let $\sigma\in{\rm Gal}_{{\rm\mathbb{Q}}}$, so $\sigma(D)\in\{D,\tau(D)\}$. If $\sigma(D)=D$, then $\sigma(y)=y$ by (8.2). Suppose $\sigma(D)=\tau(D)$. Using that $\tau(x)=-x$ and applying $\tau$ to (8.2), we see that $-x-(\tau(D)-2\infty)=\tau(\pi^{*}(y)-O_{E})=\pi^{*}(y-O_{E})$. Applying $\sigma$ to (8.2) shows that $x-(\tau(D)-2\infty)=\pi^{*}(\sigma(y)-O_{E})$. Adding the two equations from the last two sentences shows that $\sigma(y)=-y=3y$, so $\{0,y,2y,3y\}$ is indeed ${\rm Gal}_{{\rm\mathbb{Q}}}$-stable.
But $E$ is an elliptic curve with a faithful $\mu_{3}$-action, hence has CM by ${\rm\mathbb{Z}}[\omega]$, and such curves have no rational cyclic $4$-isogenies, by CM theory [8, Theorem 6.18(c)]. Contradiction! ∎
8.4. Proof of Theorem 1.1
The groups appearing in the theorem are realized infinitely often in Table 2, so it suffices to prove that $G:=P({\rm\mathbb{Q}})_{\mathrm{tors}}$ is isomorphic to one of these. Proposition 8.2 shows that the order of $G$ is of the form $2^{i}3^{j}$. Propositions 8.4 and 8.6 show that $G$ is $6$-torsion, of the form $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{m}\times({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{n}$ for some $m\geq 0$ and $n\leq 2$. Equation (7.1) and Lemma 7.2 show that $m\leq 2$ as well. Proposition 8.5 shows that if $m=2$ then $n=0$. Therefore $G$ must be one of the groups appearing in Theorem 1.1.
9. Classifying rational torsion in Pryms of ${\rm GL}_{2}$-type
Combining our knowledge of rational torsion subgroups and rational endomorphism rings of Pryms $P=P_{a,b}$, we classify the finite ${\rm End}(P)$-modules that arise as $P({\rm\mathbb{Q}})_{\mathrm{tors}}$ for Pryms $P$ of ${\rm GL}_{2}$-type.
Define the rational functions
$a_{2}(t):=\dfrac{-4(t^{2}+1)}{t^{2}(t^{2}-1)^{2}}\hskip 8.53581pt\mbox{ and }\hskip 8.53581ptb_{2}(t):=\dfrac{16}{t^{2}(t^{2}-1)^{4}}.$
Proposition 9.1.
For all but finitely many $t\in{\rm\mathbb{Q}}$, the Prym $P=P_{a_{2}(t),b_{2}(t)}$ is geometrically simple and satisfies ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{2}]$ and $({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}\subset P[{\rm\mathfrak{p}}]({\rm\mathbb{Q}})$. Conversely, if $P_{a,b}/{\rm\mathbb{Q}}$ is a Prym with these three properties then $(a,b)=(a_{2}(t)\lambda^{6},b_{2}(t)\lambda^{12})$ for some $t,\lambda\in{\rm\mathbb{Q}}$.
Proof.
First note that by Lemma 6.16, the surface $P_{a_{2}(t),b_{2}(t)}$ will be non-CM (and hence geometrically simple) for all but finitely many specializations of $t$, so for such $t$ the condition ${\rm\mathbb{Z}}[\sqrt{2}]\simeq{\rm End}(P)$ is equivalent to ${\rm\mathbb{Z}}[\sqrt{2}]\subset{\rm End}(P)$ by Corollary 6.14. By Proposition 7.14 and Corollary 6.11, $({\rm\mathbb{Z}}/3{\rm\mathbb{Z}})^{2}\subset P[\mathfrak{p}]$ and ${\rm\mathbb{Z}}[\sqrt{2}]\subset{\rm End}(P)$ if and only if $(a,b)=(-(t^{2}+1)d^{2},t^{2}d^{4})$ for some $t,d\in k$ and $16b(a^{2}-4b)\in{\rm\mathbb{Q}}^{\times 6}$. The latter is equivalent to $d=16t^{2}(t^{2}-1)\lambda^{3}$ for some $\lambda\in{\rm\mathbb{Q}}^{\times}$, proving the proposition. ∎
According to [28, Theorem 1.4], this is the largest torsion subgroup that can arise among maximal PQM abelian surfaces over ${\rm\mathbb{Q}}$ of ${\rm GL}_{2}$-type.
Note that we cannot simultaneously have ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{2}]$ and $P[3]({\rm\mathbb{Q}})\simeq{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$ since $3$ is inert in ${\rm\mathbb{Z}}[\sqrt{2}]$. However, using our family of ${\rm\mathbb{Z}}[\sqrt{6}]$-RM Pryms, we give examples of Pryms of ${\rm GL}_{2}$-type with $P[3]({\rm\mathbb{Q}})\simeq{\rm\mathbb{Z}}/3{\rm\mathbb{Z}}$. For this we define
$a_{6}(t)=36\frac{3t^{2}-1}{t^{2}(3t^{2}+1)^{2}}\hskip 8.53581pt\mbox{ and }\hskip 8.53581ptb_{6}(t)=-\frac{3888}{t^{2}(3t^{2}+1)^{4}}$
Proposition 9.2.
For all but finitely values of $t\in{\rm\mathbb{Q}}$ the surface $P_{a_{6}(t),b_{6}(t)}$ is geometrically simple and satisfies ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{6}]$ and ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\subset P[{\rm\mathfrak{p}}]({\rm\mathbb{Q}})$. Conversely, if $P_{a,b}/{\rm\mathbb{Q}}$ is a Prym with these three properties then either $P_{a,b}$ or $A_{a,b}$ is isomorphic to $P_{a_{6}(t),b_{6}(t)}$ for some $t\in{\rm\mathbb{Q}}$.
Proof.
The proof is as before, this time combining Corollary 6.12 with the two cases of Proposition 7.13. Notice that the properties of having ${\rm\mathbb{Z}}[\sqrt{6}]$-multiplication and a non-trivial rational ${\rm\mathfrak{p}}$-torsion point are preserved by duality, and the two families of Proposition 7.13 are interchanged by duality as well. Thus, it is enough to consider the first family, which leads to the functions $a_{6}(t)$ and $b_{6}(t)$. ∎
Question 2.
The minimal conductor of a geometrically simple abelian surface $A/{\rm\mathbb{Q}}$ of ${\rm GL}_{2}$-type with potential quaternionic multiplication is $3^{10}$, corresponding to a Galois orbit of weight two eigenforms $f\in S_{2}(\Gamma_{0}(243))$ with coefficients in ${\rm\mathbb{Z}}[\sqrt{6}]$ [18, Table 1]. One can show that the corresponding optimal quotient $A$ of $J_{0}(243)$ has a rational point of order $3$, a $(1,2)$-polarization, and endomorphism field $L={\rm\mathbb{Q}}(\omega)$. The same is also true for its dual. Using the results of this paper, it is not hard to show that either $A$ or its dual is isomorphic to $P_{a_{6}(t),b_{6}(t)}$ for some value of $t$. Which value of $t$ is it?
Proof of Theorem 1.5.
We have ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{D}]$ for some $D\in\{2,6\}$ by Corollary 6.14. First assume $D=2$, so that $\mathfrak{a}_{2}=(\sqrt{2})$ and $\mathfrak{a}_{3}=(3)$. Corollary 6.11 gives a parameterization (in terms of $t$ and $d$) of those $P$ with ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{2}]$. Since $A\simeq P$ in this case (again by Corollary 6.11), we have $P_{a,b}[2]({\rm\mathbb{Q}})\neq 0$ if and only if $b$ is a cube. Thus, choosing $t$ to be a cube guarantees that ${\rm\mathbb{Z}}[\sqrt{2}]/\mathfrak{a}_{2}\simeq{\rm\mathbb{Z}}/2{\rm\mathbb{Z}}\subset P[2]({\rm\mathbb{Q}})$. Proposition 9.1 gives a one-parameter family of examples with ${\rm\mathbb{Z}}[\sqrt{2}]/\mathfrak{a}_{3}\simeq{\rm\mathbb{F}}_{9}\subset P({\rm\mathbb{Q}})$. By Theorem 1.1, it remains to rule out $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}$ and ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$ as subgroups of $P({\rm\mathbb{Q}})$. If $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}\subset P({\rm\mathbb{Q}})$ then $P=P_{a,b}$ must be a specialization of the family in Corollary 7.6, and moreover $16b(w,d)(a(w,d)^{2}-4b(w,d))$ must be a sixth power. The latter is automatically a cube, so it is enough to show that it cannot be a square. Writing it out explicitly, we must show that the affine curve $Y\colon y^{2}=w(w^{3}-8)(w^{3}+1)$ has no rational points with $y\neq 0$. With the help of Magma [7] we find that the smooth projectivization $\overline{Y}$ of $Y$ is a double cover of the elliptic curve $y^{2}=x^{3}+6x-7$, which has four rational points. Checking pre-images, we find that $Y({\rm\mathbb{Q}})$ consists of the three rational points with $y=0$, as desired. Finally, we must rule out ${\rm\mathbb{Z}}/3{\rm\mathbb{Z}}\times{\rm\mathbb{Z}}/6{\rm\mathbb{Z}}\subset P({\rm\mathbb{Q}})$. If $P$ has this property, then since $P[{\rm\mathfrak{p}}]({\rm\mathbb{Q}})$ and $P_{-27}[{\rm\mathfrak{p}}]({\rm\mathbb{Q}})$ are ${\rm\mathbb{F}}_{9}$-vector spaces, one of them is isomorphic to ${\rm\mathbb{F}}_{9}$ by (7.3). Thus, exchanging $P$ with $P_{-27}$ if necessary (which is allowed since $P[2]({\rm\mathbb{Q}})\simeq P_{-27}[2]({\rm\mathbb{Q}})$), we may assume $P$ is a specialization of the family in Proposition 9.1, and hence it is enough to show that $b_{2}(t)$ is never a cube. It is then enough to show that there are no ${\rm\mathbb{Q}}$-rational points on the curve $Y^{\prime}\colon 4t^{2}(t^{2}-1)=y^{3}$ with $y\neq 0$ and $t\notin\{0,\pm 1\}$. Note the double cover $Y^{\prime}\to X$, where $X\colon 4t(t-1)=y^{3}$ is an elliptic curve (minus the origin) such that $X({\rm\mathbb{Q}})=\{(0,0),(1,0)\}$. It follows that $Y^{\prime}$ has no interesting rational points, finishing the proof in the case $D=2$.
Next consider $D=6$, so that $\mathfrak{a}_{2}=(2-\sqrt{6})$ and $\mathfrak{a}_{3}=(3+\sqrt{6})$. Corollary 6.12 then gives a parameterization (in terms of $t$ and $d$) of those $P$ with ${\rm End}(P)\simeq{\rm\mathbb{Z}}[\sqrt{6}]$. Since $A\simeq P_{-27}$ in this case, we have $P_{a,b}[2]({\rm\mathbb{Q}})\neq 0$ if and only if $b$ is a cube. Choosing $t$ so that $3t^{2}$ is a cube guarantees that ${\rm\mathbb{Z}}[\sqrt{2}]/\mathfrak{a}_{2}\simeq{\rm\mathbb{Z}}/2{\rm\mathbb{Z}}\subset P[2]({\rm\mathbb{Q}})$. Proposition 9.2 gives a one-parameter family of examples with ${\rm\mathbb{Z}}[\sqrt{2}]/\mathfrak{a}_{3}\simeq{\rm\mathbb{F}}_{3}\subset P({\rm\mathbb{Q}})$. By Theorem 1.1, it remains to rule out $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}$ and ${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$ as subgroups of $P({\rm\mathbb{Q}})$. The argument for $({\rm\mathbb{Z}}/2{\rm\mathbb{Z}})^{2}$ is exactly as before so we omit it.
It remains to rule out ${\rm\mathbb{Z}}/6{\rm\mathbb{Z}}\subset P({\rm\mathbb{Q}})$. Let $I\subset\mathcal{O}$ be the maximal two-sided ideal in $\mathcal{O}$ above $3$ (which is unique since $\mathcal{O}$ is ramified at $3$), so that $\mathcal{O}/I\simeq{\rm\mathbb{F}}_{9}$. Note that the completion $I\otimes{\rm\mathbb{Z}}_{3}\subset\mathcal{O}\otimes{\rm\mathbb{Z}}_{3}$ is generated by any element of $\mathcal{O}\otimes{\rm\mathbb{Z}}_{3}$ of minimal positive valuation. Since $\mathfrak{a}_{3}\subset{\rm\mathbb{Z}}[\sqrt{6}]$ is generated by the element $3+\sqrt{6}$ of norm $3$, and similarly $\mathfrak{p}\subset{\rm\mathbb{Z}}[\omega]$ is generated by the element $1-\zeta$ of norm $3$, we have $A[\mathfrak{a}_{3}]=A[I]=A[\mathfrak{p}]$, where $A[I]$ is the subgroup of $A$ killed by $\iota(I)$, where $\iota\colon\mathcal{O}\rightarrow{\rm End}(P_{\bar{{\rm\mathbb{Q}}}})$ is a choice of embedding satisfying the conclusion of Theorem 6.5. There is also an exact sequence
$0\to P[\mathfrak{a}_{3}]\to P[3]\to P[\mathfrak{a}_{3}]\to 0$
where the surjection is multiplication by $3+\sqrt{6}$. Thus, if $P[3]({\rm\mathbb{Q}})\neq 0$, we must also have $P[{\rm\mathfrak{p}}]({\rm\mathbb{Q}})=P[\mathfrak{a}_{3}]({\rm\mathbb{Q}})\neq 0$. It follows that $P$ is a specialization of the family in Proposition 9.2. So it suffices to show that both $b_{6}(t)$ and $16(a_{6}(t)^{2}-4b(t))$ are never cubes. These two quantities turn out to be inverses of each other modulo cubes, so it is enough to show $b_{6}(t)$ is never a cube. More precisely, it is enough to show that the curve $Y^{\prime}\colon y^{3}=12t^{2}(3t^{2}+1)$ has no rational points with $y\neq 0$. Using the double cover $Y^{\prime}\to X$ where $X\colon y^{3}=12t(3t+1)$, which is (the affine part of) an elliptic curve $\overline{X}$ with $\overline{X}\simeq{\rm\mathbb{Z}}/6{\rm\mathbb{Z}}$, we check that $Y^{\prime}({\rm\mathbb{Q}})=\{(0,0)\}$, which completes the proof. ∎